### v5

# Probabilistic Counters for Privacy Preserving Data Aggregation11footnotemark: 1

Abstract
Probabilistic counters are well-known tools often used for space-efficient set cardinality estimation. In this paper, we investigate probabilistic counters from the perspective of preserving privacy. We use the standard, rigid differential privacy notion. The intuition is that the probabilistic counters do not reveal too much information about individuals but provide only general information about the population. Therefore, they can be used safely without violating the privacy of individuals. However, it turned out, that providing a precise, formal analysis of the privacy parameters of probabilistic counters is surprisingly difficult and needs advanced techniques and a very careful approach.
We demonstrate that probabilistic counters can be used as a privacy protection mechanism without extra randomization. Namely, the inherent randomization from the protocol is sufficient for protecting privacy, even if the probabilistic counter is used multiple times. In particular, we present a specific privacy-preserving data aggregation protocol based on Morris Counter and MaxGeo Counter. Some of the presented results are devoted to counters that have not been investigated so far from the perspective of privacy protection. Another part is an improvement of previous results. We show how our results can be used to perform distributed surveys and compare the properties of counter-based solutions and a standard Laplace method.
keywords:
differential privacy, data aggregation, probabilistic counters
1 Introduction
Since Big Data related topics have been widely developed in recent years, solutions that focus on saving memory resources have become very popular. We would like to consider a standard example of such space-efficient mechanisms, namely probabilistic counters, which are used to represent the cardinality of dynamically counted events. More precisely, we would like to indicate the occurrence of $n$ events using a very small (significantly less than $\log n$) number of bits. We assume that $n$ is unknown in advance and may change. Clearly, a simple information-theoretic argument convinces us that it is not feasible if we demand an exact representation of the number of events. Nevertheless, there are some very efficient solutions that require only $\Theta(\log\log n)$ bits and guarantee sufficient accuracy for a wide range of applications. As examples, one can point most of probabilistic counters – probabilistic structures well known in the literature since the seminal Morris’ paper [1] followed by its thorough mathematical analysis by Flajolet in [2]. They are used as building blocks in many space-efficient algorithms in the field of data mining or distributed data aggregation in networks or smart metering just to mention a few ([3] or [4]).
In this paper, we investigate probabilistic counters from the privacy-protection perspective. Our analysis is based on a differential privacy notion, which is commonly considered the only state-of-the-art approach. The differential privacy has the undeniable advantage of being mathematically rigorous and formally provable, contrary to previous anonymity-derived privacy definitions. This approach to privacy-preserving protocols can be used to give a formal guarantee for privacy that is resilient to any form of post-processing. For a survey about differential privacy properties, see [5] and references therein. Analysis of protocols based on differential privacy is usually technically complex, but by using this notion, we are immune to e.g., linkage attacks (see, for example [6, 7]).
The idea behind differential privacy is as follows: for two "neighboring" scenarios that differ only in the participation of a single user, a differentially private mechanism should provide a response chosen from very similar distributions. Roughly speaking, differential privacy is described by two parameters: $\varepsilon$ – which controls a similarity of probabilities of common events – and $\delta$ – which is related to a probability of unusual events. The smaller the parameters, the better from the privacy point of view. In effect, judging by the output of the mechanism, one cannot say if a given individual (user) was taken into account for producing a given output. Intuitively, probabilistic counters should provide a high level of differential privacy since, statistically, many various numbers of events are "squeezed" into a small space of possible output results. In the case of one counter considered in our paper (MaxGeo) counter, one can find some similar, recent results about the privacy the algorithm offers. Nevertheless, the question about the value of parameters of potential differential privacy property remains open (see the discussion in Section 2) .
Moreover, when one consider small number of events $n$, an additional problem may occur, since it may be possible to distinguish that the number of events is diffrent from either $n-1$ or $n+1$ with significant probability.
In our paper, we provide a very precise analysis of two well known probabilistic counters from the perspective of preserving privacy. It turned out that this task is surprisingly complex from the mathematical point of view.
Our primary motivation is to find possibly accurate privacy parameters of two most fundamental probabilistic counter protocols, namely Morris Counter [1] and MaxGeo Counter [8]. Note that the second one is used for yet another popular algorithm — HyperLogLog [9]. One may realize that these two counters are relatively old; however, they, together with their modifications, are extensively used until these days. Morris Counter is often used in big data solutions, for instance, for measurement of network’s capabilities [10]. The most crucial examples of refinements of HyperLogLog algorithm are mentioned in Section 2.
We claim that a high-precision analysis in the case of probabilistic counters is particularly important. This is because even a mechanism with very good privacy parameters may cause a serious privacy breach when used multiple times. That is, privacy loss/information leakage acumulates over multiple releases (see e.g., [5]). Probabilistic counters in realistic scenarios may be used as fundamental primitives and subroutines in more complex protocols since the differential privacy property is immune to post-processing.
We also show that those two probabilistic counters can be used safely without any additional randomization, even in very demanding settings. It is commonly known that no deterministic algorithm can provide non-trivial differential privacy. However, Probabilistic counters possess inherent randomness, achieving desired privacy parameters. In other words, one can say that probabilistic counters are safe by design, and we do not need any additional privacy-oriented methods. In particular, what is most important, existing, working implementations do not need to be changed if we start demanding the provable privacy of a system.
Finally, we demonstrate how our results can be used for constructing a data aggregation protocol based on probabilistic counters that can be used in some specific scenarios until we want them to satisfy even more rigorous privacy properties.
To the best of our knowledge, most of the results are new and deal with protocols not considered before in the context of privacy preservation. Some other (about the so-called MaxGeo counter) improves some previous results (e.g., [11]).
1.1 Paper structure and results
Starting from this point, for the sake of clarity, we use an abbreviation DP as a shortcut for differential privacy, while this property is described by some parameters.
The main contribution of our paper is as follows:
  • 1.
    We prove that the classic Morris Counter satisfies $\left(\varepsilon(n),\delta(n)\right)$-DP with
    $\varepsilon(n)=O\left(\frac{\left(\log(n)\right)^{2}}{n}\right)$ and $\delta(n)=O\left(\max\left\{n^{-\left(\ln(n)\right)^{c-1}},n^{-1}\left(\ln(n)\right)^{-c}\right\}\right)$, for any $c>0$ (Theorem 2 in Section 6).
  • 2.
    We prove that the Morris Counter satisfies $\left(L(n),0.00033\right)$-DP property (see Definition 2), where $L(n)=-\ln\left(1-{16}/{n}\right)\approx{16}/{n}$ (Theorem 1 in Section 6). In Observation 2, we also show that the constant $16$ cannot be improved.
  • 3.
    We prove that MaxGeo Counter satisfies $\left(\varepsilon,\delta\right)$-DP property (Definition 2 is provided in Section 3) if a number of events $n$ (in Section 4 a concept of event is clarified) is at least $\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\right)\right\rceil$ (Theorem 3 in Section 4).
  • 4.
    We construct a privacy-preserving distributed survey protocol based on probabilistic counters in Section 5 and compare it with the Laplace method, which is considered as the actual state of the art of differentially private protocols and is not based on probabilistic counters.
The rest of this paper is organized as follows. First, in Section 2 we mention work related to our paper and some popular examples of other probabilistic counters, which are not considered in this paper.
In Section 3, we recall the differential privacy definition. In Section 4 we describe probabilistic counters and further, we recall the definitions of both Morris and MaxGeo Counters. Moreover, we state Fact 1, a useful reformulation of the standard definition of differential privacy for probabilistic counters. In Section 5, we demonstrate how a probabilistic counter can be used for constructing a data aggregation protocol in a very particular, yet natural, scenario. Section 6 consists of formulations of our main results for both counters. Section 7 gives some ideas about realizations of the scenario and the comparison of both counters. We also compare these solutions with the standard Laplace method (Section 7). For convenience of the reader, we provide the proofs in section 8, since they are rather technical. For more clarity, some proofs and lemmas are moved to the A. Finally, in Section 9 we present conclusions and future work propositions.
2 Previous and Related Work
In our paper, we provide a detailed analysis of some probabilistic counters from the perspective of differential privacy. Differential privacy concepts have been discussed in many papers in recent years. One can also find a well developed body of literature devoted to probabilistic counters and similar structures. Therefore, we limit the related literature review to the most relevant papers.
Differential Privacy literature
In our paper, we focus on the inherent privacy guarantees of some probabilistic structures are defined as differential privacy. The idea of differential privacy has been introduced for the first time in [12]; however, its precise formulation in the widely used form appeared for the first time in [13]. There is a long list of papers concerning differential privacy, e.g. [14, 15, 16], to mention a few.
Most of these papers focus on a centralized (global) model, namely a database with a trusted party holding it. See that in our paper, despite the distributed setting, we have the same (non-local) trust model. In particular we assume an existence of a curator that is entitled to gather and see all participants’ data in the clear and release the computed data to a wider (possibly untrusted) audience. Comprehensive information concerning differential privacy can be found in [5].
Probabilistic counters and their applications
The idea of probabilistic counters, along with the well-known Morris Counter was presented in the seminal paper [1]. The aim was to construct a very small data structure for representing a huge set or events of some kind. In our paper, we concentrate on the Morris Counter analyzed in detail in [2]. The second structure discussed in our paper is MaxGeo Counter introduced and analyzed in [8]. More detailed and precise analysis can be found in [17]. The most important application of MaxGeo Counter can be found in [9], where the authors propose the well-known HyperLogLog algorithm. Its practical applications are widely described in [18]. There are several widely used improvements of HyperLogLog algorithm: HyperLogLog+ [18], Streaming HyperLogLog with sketches based either on historic inverse probability [19] or martingal estimator [20] or empirically adjusted HyperBitBit (proposed by R. Sedgewick [21]). The main goal of these adjustments is a reduction of the memory requirements (see e.g., [22] or [23]). For instance, some of the above solutions are used in database systems for queries’ optimization or for document classification purposes. Moreover, MaxGeo counter was used in [24], for an adjustment of ANF tool, developed for data mining from extensive graphs, which enables it to answer many different questions based on some neighborhood function defined on the graph.
Unsurprisingly one of the main applications of the approximate counter is to compute the size of a database or its specific subset. A set of such applications can be found in [25]. In [26], the authors use Morris Counter for online, probabilistic, and space-efficient counting over streams of fixed, finite length. Authors of [27] proposed an application of a system of Morris Counters for flash memory devices. Another application, presented in [28], is a revisited version of Morris Counter designed for binary floating-point numbers. In [29], Morris Counter is used in a well-known problem of counting the frequency moments of long data streams. Authors of [30] focused on making probabilistic counters scalable and accurate in concurrent settings. Paper on probabilistic counters in hardware can be found in [31]. A slightly modified version of Morris Counter called Morris+ was recently introduced in [32] with the proof of its optimality in terms of accuracy–memory trade-off.
In random graphs theory, Morris Counter is usually connected to greedy structures. For instance, in an arrangement of a random labeled graph in Gilbert model $G(n,p)$, it is possible to construct a greedy stable set $S_{n}$, which size has the same distribution as Morris Counter $M_{n}$ of the base $a=(1-p)^{-1}$ (see e.g., [33] or [34] for fundamentals of random graph theory).
There are many other birth processes that are quite similar to Morris Counter, which are applicable in a variety of disciplines like biology, physics, or the theory of random graphs. Short descriptions of such examples can be found in [35].
When talking about probabilistic counters, it is worth to mention about Bloom filters [36], which are space-efficient probabilistic data structures that are the representations of sets. There exists a probabilistic counter which approximates the number of elements represented by the given Bloom filter [37].
Other common examples of probabilistic counters are $F_{p}$ counters [38, 39], which approximate the $p$-th moments of frequencies of occurrences of different elements in the database.
Let us also mention a paper [40] where one can find numerous applications of similar constructions for creating pseudorandom sketches in Big Data algorithms.
Notice that the variety of possible applications of probabilistic counters creates an opportunity to exploit inherent differential privacy properties. However, a new challenge arises — to calculate the parameters of differential privacy for those counters, which are not connected straightforwardly with Morris or MaxGeo Counters.
Probabilistic counters and preserving privacy
Some probabilistic counters and similar structures were previously considered in terms of privacy preservation. We mention only the papers strictly related to the algorithms discussed in our paper (i.e., Morris Counter and MaxGeo).
The authors of [41] show that in the scenario of using different types of probabilistic counters for set cardinality estimation with the Adversary being able to extract the intermediate values of the counter, the privacy is not preserved. Note that, in this paper, we perform data aggregation instead of cardinality estimation. Moreover, we assume the Adversary is not able to extract any intermediate values from the counter. That is, we consider a global model, while the result from [41] assumes the settings closer to the classic local model [5].
One of the main results of this submission is a careful and tight analysis of Morris Counter from the context of privacy-preserving. To the best of our knowledge such analisys has not been provided so far. Our second contribution is an analogous analysis of the MaxGeo.
There are a few very recent papers presenting privacy-preserving protocols that use Flajolet—Martin sketch as a building block. In [25], so called Flajolet—Martin sketch was introduced. In [11] authors consider a general concept of a probabilistic counter, based on several MaxGeo counters and its differential privacy. Nevertheless, they incorrectly call it a generalization of Flajolet—Martin sketch (they probably confuse Flajolet—Martin sketch with LogLog sketch). We concern a concept of MaxGeo counter, which is a core of LogLog or HyperLogLog sketch, however, it can be used in other arrangements as well. These papers in some cases provide an analysis of the privacy guaranteed by Flajolet—Martin with the global model. In all the cases, the conclusion is positive in the sense that the protocol itself provides some level of differential privacy without adding extra randomness.
Beneath papers provide an analysis of privacy guaranteed by the sketches related to LogLog algorithm, which can be seen as the processing of the fundamental MaxGeo counters. From this point of view, our contribution about MaxGeo counters has a larger applicative potential.
  • 1.
    In [11] authors consider, among others, a sketch that can be seen as a particular application of MaxGeo counter. They introduce its differentially private version via trick (adding artificial utilities) and provide its accuracy when used to count the number of elements in multisets.
    Accidentally, a proof of basic theorem from [11] uses an incorrect argument (inappropriate utilization of Hoeffding’s inequality), so it is difficult to compare the results precisely. Nevertheless, the overlap of results between our paper and [11] is only partial.
  • 2.
    In [42], authors consider LogLog sketch as a subroutine. After a careful analysis, they show that it is asymptotically $(\varepsilon,\delta=\mathrm{negl}(\lambda))$-DP (with respect to the numbers of different elements), when the number of elements counted by the mechanism is at least $8K\lambda\max(\frac{1}{\varepsilon},1)$, where $K$ is some accuracy parameter, $\lambda$ is some security parameter and $\mathrm{negl}(x)$ is some negligible function of argument $x$ (Theorem 4.2 in [42]). Nevertheless, the analysis does not explain how to choose parameters $K$ and $\lambda$ in order to obtain $(\varepsilon,\delta)$-DP for a given $\varepsilon$ and $\delta$ parameters. Moreover, a consideration of asymptotic behavior (with respect to the number of unique elements $n$) is not relevant when the hash function restricts the possible result to the size bounded by its domain. Our analysis of MaxGeo counter provides exact (non-asymptotic) dependence between $n$ and parameters $\varepsilon$ and $\delta$.
Let us also mention that some other pseudorandom structures have been analyzed from the perspective of differential privacy. For example, in [43], the authors considered Bloom filters as a means for constructing privacy-preserving aggregation protocol.
3 Differential Privacy Preliminaries
In this section, we briefly recall differential privacy. For more details see, e.g. [5]. We denote the set of (positive) natural numbers by $\mathbb{N}$ and the set of all integers by $\mathbb{Z}$. Moreover, let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $a,b\in\mathbb{Z}$ let us define a discrete interval $[a,b]\cap\mathbb{Z}$ by $[a:b]$. We also define $[n]=\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$. We assume that there exists a trusted curator who holds, or securely obtains, the data of individuals in a (possibly distributed) database $x$. Every row of $x$ consists of the data of some individual. By $\mathcal{X}$, we denote the space of all possible rows. The goal is to protect the data of every single individual, even if all users except one collude with an adversary to breach the privacy of this single, uncorrupted user. On the other hand, the curator is responsible for producing a release – a possibly accurate response to a requested query. This response is then released to the public, which is allowed to perform statistical analysis on it. The differential privacy is, by design, resilient to post-processing attacks, so even if the adversary obtains the public release, he will not be able to infer anything about specific individuals participating in this release.
For simplicity, we interpret databases as their histograms in $\mathbb{N}_{0}^{|\mathcal{X}|}$, so we may just focus on unique rows and the numbers of their occurrences.
Definition 1 (Distance between databases)
The $\ell_{1}$ distance between two databases $x,y\in\mathbb{N}_{0}^{|\mathcal{X}|}$ is defined as
$\|x-y\|_{1}=\sum_{i\in\mathcal{X}}|x_{i}-y_{i}|,$
where $x_{i}$ and $y_{i}$ denote the numbers of occurrences of an item (an individual) $i$ in the databases $x$ and $y$ respectively.
One can easily see that $\|x-y\|_{1}$ measures how many records differ between $x$ and $y$. Moreover, $\|x\|_{1}$ measures the size of the database $x$.
A privacy mechanism is a randomized algorithm used by the curator that takes a database as input and produces the output (the release) using randomization.
Definition 2 (Differential Privacy (formulation from [5]))
A randomized algorithm $\mathcal{M}$ with domain $\mathbb{N}^{|\mathcal{X}|}$ is ($\varepsilon,\delta$)-differentially private (or ($\varepsilon,\delta$)-DP), if for all $\mathcal{S}\subseteq$ $\mathcal{M}$) and for all $x,y\in\mathbb{N}^{|\mathcal{X}|}$ such that $\left\|x-y\right\|_{1}\leqslant 1$ the following condition is satisfied:
$\mathbb{P}\left(\mathcal{M}(x)\in\mathcal{S}\right)\leqslant\mathrm{exp}(\varepsilon)\cdot\mathbb{P}\left(\mathcal{M}(y)\in\mathcal{S}\right)+\delta,$
where the probability space is over the outcomes of the mechanism $\mathcal{M}$.
When $\delta=0$, $\mathcal{M}$ is called ($\varepsilon$)-DP mechanism.
An intuition of $(\varepsilon,\delta)$-DP is as follows: if we choose two consecutive databases (that differ exactly on one record), the mechanism will likely return indistinguishable values. In other words, it preserves privacy with high probability, but it is admissible for a mechanism to be out of control with negligible probability $\delta$.
Example 1
(Laplace noise) In the central model, a standard, widely used mechanism with $(\varepsilon)$-DP property is the so-called Laplace noise. A variable $X$ has Laplace distribution with parameter $\lambda$ (denoted as $X\sim\mathcal{L}(\lambda)$), if its probability density function is
$f(x)=\frac{1}{2\lambda}\exp\left(-\frac{|x|}{\lambda}\right)~{}.$
Then $\mathbb{E}\left(X\right)=0$ and $\mathrm{Var}(X)=2\lambda^{2}$.
Let $c(x)$ be the number of rows in $x$, that satisfy a given property. Note that $c$ in the differential privacy literature is usually referred to as count query. Imagine that an aggregating mechanism is defined as follows: $\mathcal{M}(x)=c(x)+\mathcal{L}(\varepsilon^{-1})$. Then $\mathcal{M}$ is $(\varepsilon)$-DP (for more precise properties of Laplace noise, see [5]).
4 Probabilistic Counters — preliminaries
This paper focuses on probabilistic counters, denoted further by $M$. The notion of the probabilistic counter is ambiguous in literature. It is a stochastic process that can be interpreted as a mechanism defined on the space of all possible inputs, which should estimate some goal value in some sense. The exact definition of this approximation is not crucial from DP-point of view, thence we do not consider it in this paper.
Each increase of the data source being counted by the probabilistic counter is called an increment request. Due to the randomized nature of probabilistic counters, each may change the value of the counter, but not necessarily. We will further indicate the single increment request by ${}^{\prime}1^{\prime}$. For the sake of generality, we also assume that the counter can get as an input ${}^{\prime}0^{\prime}$, and in such case, it simply does nothing. This is useful for real-life scenarios, e.g., data aggregation (see Section 5). Obviously, only increment requests impact the counter’s final result; hence, we indicate the counter’s value after $n$ increment requests by $M_{n}$, and we are not considering the number of the rest of the rows.
In Figure 1, one can see a graphical representation of the probabilistic counter. As mentioned, increment requests are indicated by ${}^{\prime}1^{\prime}$ and other rows by ${}^{\prime}0^{\prime}$ input. The dice depicts the randomness. The $\mathbf{X}$-mark stands for no action.
Figure 1: Graphical depiction of the probabilistic counter.
We emphasize that the probabilistic counter depends on the number of increment requests. We want to show that if we reveal its final value, then it does not expose any sensitive data about any single record. Moreover, note that if $x$ and $y$ differ only by one ${}^{\prime}0^{\prime}$ input, then $\mathbb{P}\left(M(x)\in\mathcal{S}\right)=\mathbb{P}\left(M_{n}\in\mathcal{S}\right)=\mathbb{P}\left(M(y)\in\mathcal{S}\right)$, where $n$ is the number of increment requests both for $x$ and $y$. See that then the condition in Definition 2 is trivially fulfilled. Hence for our convenience, in this paper, we use to mark only the number $n$ of increment requests provided by individuals when talking about the probabilistic counter $M_{n}$.
Fact 1
Let $M$ be a probabilistic counter with a discrete $Range(M)=A$. Moreover assume that for all $n,m\geqslant 1$, such that $|n-m|\leqslant 1$, there exists such $S_{n}\subset A$ that for all $s\in S_{n}$
$\mathbb{P}\left(M_{n}=s\right)\leqslant\mathrm{exp}(\varepsilon)\cdot\mathbb{P}\left(M_{m}=s\right)~{}$ (1)
and
$\mathbb{P}\left(M_{n}\notin S_{n}\right)\leqslant\delta~{}.$ (2)
Then $M$ is $(\varepsilon,\delta)$-DP.
Note that, for our setting, Fact 1 is fully compatible with the intuition of regular differential privacy (Definition 2). Indeed, Fact 1 can be easily derived from the observation that any set $B\subset A$ is a disjoint union of $B\cap S_{n}$ and $B\cap S_{n}^{\prime}$.
Remark that $\varepsilon$ and $\delta$ in Fact 1 can be also treated as functions of a parameter $n$, i.e. with respect to the number of increment requests. Thence, we can consider either differential privacy of $M_{n}$ variable, when $n$ is known or ($\varepsilon(n),\delta(n))$-DP of probabilistic counter $M$. The second variant let us provide a precise dependence of privacy parameters of the counter as the number of increment requests gets large.
4.1 Morris Counter
We begin with a short description of Morris Counter (originally referred to as an approximate counter [1, 2]). Let us fix $a>1$. Algorithm 1 is a very simple pseudocode of Morris Counter [1].
$M\leftarrow 1$;
2 while receive request do
3       generate $r\sim\mathrm{Uni}([0,1])$;
4       if $r<a^{-M}$ then
5             $M\leftarrow M+1$;
6
Algorithm 1 Morris Counter Algorithm
Roughly speaking, we start with $M=1$. Each incoming increment request triggers a random event. This event increments the counter ($M\leftarrow M+1$) with probability $a^{-M}$ ($r\sim\mathrm{Uni}([0,1])$ generates a number uniformly at random from the interval $[0,1]$, using in practice some Pseudo Random Number Generator). Note that this approximate counting protocol can be easily distributed. Indeed, any entity who wants to increment the counter only has to send the request to increment it. These requests can be queued on the server and resolved one after another. A detailed description of the approximate counting method can be found in [1, 2]. Throughout this article, we examine only a standard Morris Counter i.e., with the base $a=2$. Morris Counter can also be defined recursively.
Definition 3
Morris Counter is a Markov process $(M_{n},n\in\mathbb{N}_{0})$ which satisfies:
$\mathbb{P}\left(M_{0}=1\right)=1~{},$
$\mathbb{P}\left(M_{n+1}=l|M_{n}=l\right)=1-2^{-l}~{},$
$\mathbb{P}\left(M_{n+1}=l+1|M_{n}=l\right)=2^{-l}~{},$
for any $l\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$.
Note that Definition 3 can be derived directly from a run of Algorithm 1. From now on let $\mathbb{P}\left(M_{n}=l\right)=p_{n,l}$. Directly from Definition 3 we get the following recursion:
$p_{n+1,l}=(1-2^{-l})p_{n,l}+2^{-l+1}p_{n,l-1}~{}$ (3)
for $l\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ with starting and boundary conditions $p_{0,1}=1$, $p_{0,l}=0$ for $l\geqslant 2$ and $p_{n,0}=0$ for $n\in\mathbb{N}_{0}$.
Accuracy versus Differential Privacy
The accuracy of Morris Counter has been thoroughly analyzed in various classical papers. The first detailed analysis was proposed by Ph. Flajolet in [2]. In this part we present the essence of theorems presented in this paper, which will be useful later on.
First, we provide the asymptotics of the expected value and the variance of Morris Counter, with precise numerical approximations of constants:
Fact 2
Let $M_{n}$ denote Morris Counter after $n$ successive increment requests. Then this random variable has an expected value $\mathbb{E}\left(M_{n}\right)\approx\log(n)-0.27395$ (in this paper $\log$ states for binary logarithm) and a variance $\mathrm{Var}(M_{n})\approx 0.763014$.
Realize that Fact 2 guarantees high concentration of $M_{n}$ around its average — a characteristic desirable in order to satisfy differential privacy definition. Fact 2 also justifies a definition of moving discrete intervals:
$I_{n}=[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4]\cap[n+1]~{},$ (4)
which will emerge as a crucial point of our further considerations of this Markov process in terms of differential privacy in this section. Let us mention, that usually $[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4]\subseteq[n+1]$, so one may think that $I_{n}$ are symmetric discrete intervals of length $8$, centered at $\lceil\log(n)\rceil$.
The lion’s share of applications of Morris Counter is based on counting a number of occurrences, that is, the number of increment requests. In order to estimate this value, we may use  (3) and simply obtain $\mathbb{E}\left(2^{M_{n+1}}\right)=\mathbb{E}\left(2^{M_{n}}\right)+1$, so together with the assumption $M_{0}=1$ we obtain
$\mathbb{E}\left(2^{M_{n}}\right)=n+2~{}.$ (5)
Hence $22^{M_{n}}-$ 2 is an unbiased estimator of the number of increments $n$. Remark that $n$ can be saved in $\lceil\log(n)\rceil$ bits. On the other hand, Fact 2 shows that on average, $\log(\log(n))+O(1)$ bits are required to store $M_{n}$. As announced earlier, this is the crucial advantage of Morris Counter. Moreover, analogously to (5) we may obtain
$\mathrm{Var}(2^{M_{n}}-2)=\frac{n(n+1)}{2}~{}.$ (6)
Formulas (5) and (6) will be used in the analysis of data aggregation example in Section 5.
4.2 MaxGeo Counter
We begin with a short description of MaxGeo Counter. Algorithm 2 shows its pseudocode. Informally, for each increment request, the server has to generate a random variable from the geometric distribution $\mathrm{Geo}({1}/{2})$ (ranged in $\mathbb{N}$). The final result is the maximum taken over all these generated random variables.
$C\leftarrow 1$;
2 while receive request do
3       generate $r\sim\mathrm{Geo}({1}/{2})$;
4       $C\leftarrow\max\{C,r\}$;
5
6return $C$;
Algorithm 2 MaxGeo Counter Algorithm
The expectation and variance of the maximum of $n$ i.i.d. geometric variables have already been analyzed in the literature. For instance, Szpankowski and Rego [8] provided exact formulas for such variables’ expected value and variance. However, they are impractical for applications for large $n$. Hence they also provided asymptotics (here, for a maximum of $n$ independent $\mathrm{Geo}(1/2)$ distributions): $\mathbb{E}\left(M_{n}\right)=\log(n)+O(1)$ and $\mathrm{Var}(M_{n})=\log(n)+O(1)$ and thus, similarly to Morris Counter, there are only $\log(\log(n))+O(1)$ bits required on average to save MaxGeo Counter after $n$ increment requests.
4.3 General Probabilistic Counting with Stochastic Averaging
Here we recall briefly a General Probabilistic Counting with Stochastic Averaging algorithm, based on the original idea from [25]. Assume that there are $m$, initially empty lots related to some independent copies of some probabilistic counter. For each increment request, we connect it with one of the groups uniformly at random. Finally, we perform incrementation requests separately and independently for each lot, obtaining $M[1],M[2],\ldots,M[m]$.
Without delving into details, for original PCSA algorithm, $\mathbb{E}\left(M_{n}\right)\approx\log(\varphi n)$, where $M_{n}$ is a value of a specific probabilistic counter connected with PCSA after $n$ increment requests and $\varphi$ is some magic constant. If we denote the mean of these $m$ counters after the total number of increment requests $n$ by $\sigma_{n}(m)$, then we may introduce the statistic:
$\Xi_{n}(m)=\left\lfloor\frac{m}{\varphi}2^{\sigma_{n}(m)}\right\rfloor~{}.$
Then (according to [25]), for any $m=2^{k}$, $k\in\mathbb{N}$, $\mathbb{E}\left(\Xi_{n}(m)\right)\approx n\left(1+\frac{0.31}{m}\right)$ and $\mathrm{Var}(\Xi_{n}(m))=n^{2}\left(\frac{0.61}{m}\right).$
Note that averaging reduces the variance of the probabilistic counter. Remark that ”Stochastic Averaging” in PCSA algorithm refers to the random choice of the number of entities in each group, and it slightly differs from the standard averaging solution via the Monte Carlo method with groups of equal size.
An important conclusion is that we may apply the idea of original PCSA in general, for any probabilistic counters.
4.4 LogLog counter
In [44] a LogLog algorithm was proposed. It is based on $m=2^{k}$ counters $(M[j])_{j=1}^{m}$, where $k>0$. We may interpret this algorithm in the context of probabilistic counters. In such a scenario, it takes a hashed value (binary sequence) as an input on every increment request. The first $k$ bits of the hash determine which of the $m$ counters should be incremented (the index $j$ is chosen as the decimal representation of the sequence restricted to these first $k$ bits; hence it translates the increment request of the LogLog counter to the increment request of one of the $m$ internal counters). Consider the first non-zero bit of a tail of the sequence (starting from $(k+1)$-th bit). Its position $R$ in this tail follows the $\mathrm{Geo}\left(\frac{1}{2}\right)$ distribution assuming the uniform distribution of the input sequences. If $R>M[j]$, then $M[j]$ should become $R$. Otherwise it does not change.
Therefore, LogLog counter is, in fact, a general PCSA that uses $m$ MaxGeo counters. It can be used to estimate the cardinality of the increment requests $n$ using the following estimator:
$\mathrm{LogLog}_{n}^{(m)}=\alpha_{m}m2^{\frac{1}{m}\sum\limits_{j=1}^{m}M_{n}[j]}~{},$
with the scaling constant given by the formula
$\alpha_{m}=\left(\Gamma\left(-\frac{1}{m}\right)\frac{2^{-1/m}-1}{\ln(2)}\right)^{-m}~{},$
where $\Gamma$ is Euler’s Gamma function.
It is worth mentioning that $\alpha_{m}$ is an increasing sequence and $\alpha_{m}\approx 0.79$ for $m\geqslant~{}64$ (e.g. $\alpha_{8}\approx 0.69763\ldots$, $\alpha_{64}=0.78356\ldots$ and the limit is $\alpha_{\infty}=0.79402\ldots$).
We are interested in the expectation and accuracy of the cardinality estimator, which can be briefly described as follows: $\mathbb{E}\left(\mathrm{LogLog}_{n}^{(m)}\right)\approx n$ and $\mathrm{Var}(\mathrm{LogLog}_{n}^{(m)})\approx\frac{1.69n^{2}}{m}$ (check [44] for more details). However, in order to control the LogLog counter, on average about $m\log\log\left(\frac{n}{m}\right)$ bits of memory are needed.
An interesting fact is that the expectation of single MaxGeo counter is logarithmic and the estimator of $n$ is of the form $Cm2^{\sigma_{n}(m)}$, for some constant $C$ (just like in the case of original PCSA).
Since LogLog counter is an effect of processing of some MaxGeo counters, therefore its differential privacy is based on the same property of the auxiliary counters.
4.5 HyperLogLog
The maximum of geometric variables is also used as a primitive in well known HyperLogLog algorithm (see [9]). Therefore its privacy properties are important both from the theoretical and practical point of view. Essentially, in HyperLogLog we perform the general PCSA algorithm, but the final estimation is somehow different:
$\mathrm{HyperLogLog}_{n}^{(m)}:=\alpha_{m}m^{2}\left(\sum_{j=1}^{m}2^{-M_{n}[j]}\right)^{-1},$
where $\alpha_{k}$ is a constant dependent only on $k$ (see [9] for more details). It is worth noting that HyperLogLog related algorithms (mentioned in Section 1) are the best-known procedures designated for cardinality estimation, and they are close to optimum [45]. According to [9], for $m=2^{k}$, where $k\geqslant 4$,
$\mathbb{E}\left(\mathrm{HyperLogLog}_{n}^{(m)}\right)=n(1+\psi_{3}(n)+o(1)),\text{ with }|\psi_{3}(n)|<5\cdot 10^{-5}$
and
$\mathrm{Var(HyperLogLog)_{n}^{(m)}}=n^{2}\left(\frac{\beta_{m}}{\sqrt{m}}+\psi_{4}(n)+o(1)\right)^{2},\text{ with }|\psi_{4}(n)|<5\cdot 10^{-4}~{},$
where $\beta_{m}\stackrel{{\scriptstyle m\rightarrow\infty}}{{\longrightarrow}}\sqrt{2\log(2)-1}=1.03896\ldots$ and $\beta_{m}\leqslant 1.106$ for $m\geqslant 16$.
5 Privacy-Preserving Survey via Probabilistic Counters
In this section, we present an example scenario for data aggregation using probabilistic counters. We assume there is a server (alternatively, we call it aggregator) and a collection of nodes (e.g., mobile phone users), and we want to perform a boolean survey with a sensitive question. Namely, each user sends ${}^{\prime}0^{\prime}$ if his/her answer is no and ${}^{\prime}1^{\prime}$ if the answer is yes. We assume that the connections between users and the server are perfectly secure and the data can safely get to the trusted server. This can be performed using standard cryptographic solutions. The goal of the server is to publish the sum of all ${}^{\prime}1^{\prime}$ responses in a privacy-preserving way. Such a goal could obviously be achieved by simply collecting all the data and adding an appropriately calibrated, Laplace noise (see [5]). However, we aim to show that probabilistic counters have inherently sufficient randomness to be differentially private without any auxiliary randomizing mechanism.
We can present the general scenario in the following way:
  1. 1.
    each user sends his/her bit of data to the server using secure channels,
  2. 2.
    server plugs the data points sequentially into the counter,
  3. 3.
    if the data point is ${}^{\prime}1^{\prime}$, the counter receives increment request, otherwise, the data is ignored,
  4. 4.
    each increment request is being processed by the counter and may lead (depending on randomness) to an increase of the value of the counter,
  5. 5.
    when all data is processed, the value of the counter is released to the public.
Note that we assume that the Adversary has access only to the released value. We also released just the counter value itself, which does not estimate ${}^{\prime}1^{\prime}$ responses. Such estimation is a function of released value, which is different for Morris or MaxGeo Counter. There can also be various ways to estimate the actual number using counter value. However, this does not matter for our case, as differential privacy is, conveniently, fully resilient to post-processing (see [5]). The graphical depiction of our considered scenario is presented in Figure 2.
Figure 2: Scenario for data aggregation using probabilistic counters. We assume that the Adversary does not have any way to extract information from within the rectangle.
Adversary
Our assumptions about the Adversary are the same as in most differential privacy papers. Namely, he may collude with any subset of the participants (e.g., all except the single user whose privacy he wants to breach). On the other hand, the aggregator is trusted. See that even though we have a distributed system in mind, this is, in fact, a central differential privacy scenario. We do not assume pan-privacy. It means that the algorithm’s internal state is not subject to the constraints of differential privacy. Obviously, if the Adversary would know the internal state of the counter at any time or could observe whether, after receiving data from a specific user, the server has to perform computations to potentially increment the counter (implying a ${}^{\prime}1^{\prime}$ response) or not, he would easily violate the privacy. We also do not assume privacy under continual observation. The survey is not published iteratively but one time only after it is finished. To sum it up, the Adversary cannot
  • 1.
    extract or tamper with the internal state of the counter,
  • 2.
    extracts any information from the server or channels between users and the server.
The Adversary can
  • 1.
    collude with any subset $C$ of the participants (e.g. know their data or make them all send ${}^{\prime}0^{\prime}$ to the server (intentially)) in order to breach the privacy of user not belonging to $C$,
  • 2.
    obtain the final result of the aggregation and perform any desired post-processing on it.
6 Formulation of main results
In this section we state main results. Proofs are postponed to Section 8 for the convenience of the reader.
6.1 Morris Counter Privacy
In this subsection we investigate Morris Counter in terms of $(\varepsilon,\delta)$-DP. Here we present the main result:
Theorem 1
Let $M$ denote the Morris Counter and assume $|n-m|\leqslant 1$. Then
$\mathbb{P}\left(M_{n}=l\right)\leqslant\left(1-\frac{16}{n}\right)^{-1}\cdot\mathbb{P}\left(M_{m}=l\right)+\delta,$
where $\delta<0.00033$, so $M$ is $\left(L(n),0.00033\right)$-DP with
$L(n)=-\ln\left(1-\frac{16}{n}\right)=\frac{16}{n}+\frac{128}{n^{2}}+O(n^{-3})\leqslant\frac{16}{n-8}~{}.$
An explanation is postponed to subsection 8.1.
6.2 General result on Morris’ Counter privacy
In this part, we show that the Morris’ Counter guarantees privacy with both parameters tending fast to zero. The analysis is based on the observations from the previous case. However, instead of $I_{n}$ (see Section 4.1 for the discussion), we consider intervals
$J_{n}(c)=[\lceil\log(n)\rceil-\lceil c\log(\ln(n))\rceil:\lceil\log(n)\rceil+\lceil c\log(\ln(n))\rceil]\cap[n+1]~{},$
where $c$ is some positive constant such that $\lceil c\log(\ln(n))\rceil\geqslant 1$, for big enough $n$.
We can now state our next contribution:
Theorem 2
Let $M$ denote the Morris Counter. If $c>0$ satisfies $\lceil c\log(\ln(n))\rceil\geqslant 1$, then $M$ is $\left(\varepsilon(n),\delta(n)\right)$-DP with parameters $\varepsilon(n)=O\left(\frac{\left(\log(n)\right)^{2}}{n}\right)$ and
$\delta(n)=O\left(n^{-\left(\ln(n)\right)^{c-1}}+n^{-1}\left(\ln(n)\right)^{-c}\right)$.
A proof is given in subsection 8.2
6.3 MaxGeo Counter Privacy
In this subsection, we present a theorem that shows the privacy guarantees of MaxGeo Counter. Assume that we have $n$ increment requests. In the case of MaxGeo Counter it means that we generate random variables $X_{1},\ldots,X_{n}$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. Ultimately the result of the counter is maximum over all $X_{i}$’s, namely $X=\max(X_{1},\ldots,X_{n})$. Now we are ready to present our second main contribution.
Theorem 3
Let $M$ denote the MaxGeo Counter, and $n$ denote the number of increment requests. Consider $m$ such that $|n-m|\leqslant 1$. Fix $\varepsilon>0$ and $\delta\in(0,1)$ and let
$l_{\varepsilon}=\left\lceil\log\left(\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\right)\right\rceil~{}.$
If
$n\geqslant\frac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}\left(\approx-\frac{\ln(\delta)}{\varepsilon}\right)~{},$ (7)
then
$\mathbb{P}\left(M_{n}\in S\right)\leqslant\mathrm{exp}(\varepsilon)\cdot\mathbb{P}\left(M_{m}\in S\right)+\delta,$
so $M$ is $(\varepsilon,\delta)$-DP.
A proof is postponed to subsection 8.3.
7 Practical applications
7.1 Discussion on arragements of the survey scenario
We usually consider probabilistic counters to be some kind of estimator. There is always a trade-off between its precision and the quality of privacy. Ultimately, it means that for fixed privacy parameters $(\varepsilon,\delta)$ we can calculate the minimum number of increment requests necessary to satisfy given privacy parameters. This can be done by artificially adding them before actually collecting data. Of course, it has to be taken into account that the initially added value should be subtracted from the final estimation of the appropriate cardinality before publication, and this change can impact the precision of the estimation (especially when the expected number of increment requests is very small). If we can perform such a preprocessing, then for every $(\varepsilon,\delta)$, we can easily know how many artificial counts have to be added. Nevertheless, we have to be aware that if the total number of increment requests is small, we may obtain poor approximation, so this approach should be used whenever the privacy is much more important than the precision.
Note that, in light of our theorems 1 and 3, both Morris Counter and MaxGeo Counter preserve differential privacy in such a scenario. Assume at least $n$ users holding ${}^{\prime}1^{\prime}$, therefore at least $n$ incrementation requests. See that we can either know it based on domain knowledge (e.g., we expect that at least some fraction of users will send ${}^{\prime}1^{\prime}$ based on similar surveys) or add $x$ counts to the counter artificially initially. The number $x$ should be chosen according to the maximal amenable value of $\varepsilon$ parameter for a given application, but we recommend choosing rather small values of $x$. Obviously, in the case of artificial counts, it has to be taken into account when estimating the final sum. Using Morris Counter we obtain $\left(L(n),0.00033\right)$-DP with
$L(n)=-\ln\left(1-\frac{16}{n}\right)\leqslant\frac{16}{n-8}~{}.$
For instance,
Example 2
Consider a result of Morris Counter with a small number $n$ of increment requests (for example, a number of respondents suffering from a rare illness). Therefore, we will likely require the $\varepsilon$ parameter to be at most some threshold, e.g. $1$. Therefore, from Theorem 1, we should add $x$ counts where $L(x)\leqslant{16}/{(x-8)}\leqslant 1$, so $x\geqslant 24$. Note that we do not include $n$ in the above formula since it is not known in advance. Therefore, using Morris Counter, the above survey is at least $(L(n+24),0.00033)$-DP. However, the estimator should be modified as well, i.e., $M^{\prime}_{n}=M_{n+24}$, so $\hat{n^{\prime}}=\max\{2^{M^{\prime}_{n}}-26,0\}$ (since $2^{M^{\prime}_{n}}$ may be smaller than $26$).
On the other hand, using MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$ (see Theorem 3).
Example 3
Assume that we have at least $n=200$ increment requests. From Theorem 1, we have $L(n)\leqslant{16}/{(n-8)}\leqslant 0.08334$. Therefore, using Morris Counter, the above survey is $(0.08334,0.00033)$-DP.
On the other hand, using MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$. Here we present an example.
Example 4
Let $\varepsilon=0.5$ and $\delta={1}/{D^{2}}$, where $D$ is the number of all survey participants. After using our theorem and straightforward calculations, we have $n\geqslant 7\ln(D)~{}.$ Say we will have $\lfloor\exp(20)\rfloor$ participants. Then if we have at least $140$ increment requests, we satisfy $(0.5,{1}/{D^{2}})$-DP.
Note that from a differential privacy perspective, both the general PCSA algorithm and HyperLogLog can be seen as arbitrary postprocessing performed on the $m$ MaxGeo counters. Moreover, as each response goes to one counter only, they are independent of each other, so we can use the parallel composition theorem (see [5]).
Observation 1
Assume we have $k$ MaxGeo Counters $M[1],\ldots,M[m]$, which are used either in HyperLogLog or PCSA algorithm. If $j$th MaxGeo Counter is $(\varepsilon_{j},\delta_{j})$-DP then the chosen algorithm is $(\max\limits_{i}\varepsilon_{i},\max\limits_{i}\delta_{i})$-DP.
7.2 Comparison of Morris and MaxGeo Counters
In this subsection, we compare data aggregation algorithms’ privacy and storage properties based on one of the investigated counters or the standard Laplace method.
We start with auxiliary remarks for the privacy of MaxGeo Counter. For instance, see that if $\delta$ and $n$ are fixed, then from Theorem 3 and $l_{\varepsilon}\leqslant\left\lceil\ln\left(1+\varepsilon^{-1}\right)\right\rceil$ we obtain that
$\varepsilon(n)\geqslant\left(2^{\left\lfloor-\log\left(1-\delta^{\frac{1}{n}}\right)\right\rfloor}-1\right)^{-1}=:\varepsilon_{0}(n)~{}.$ (8)
We want to optimize $\varepsilon(n)$, so we are going to consider $\varepsilon_{0}(n)$ defined as the right-hand side of (8). In order to limit it let us consider the following function of $x\in\mathbf{R}_{+}$:
$\psi(x,\delta):=\left(\left(1-\delta^{\frac{1}{x}}\right)^{-1}-1\right)^{-1}=-\frac{\ln(\delta)}{x}+\frac{\ln(\delta)^{2}}{2x^{2}}-\frac{\ln(\delta)^{3}}{6x^{3}}+O(x^{-4})~{}.$ (9)
Naturally, then $\varepsilon_{0}(n)\geqslant\psi(n,\delta)=-{\ln(\delta)}/{n}+O(n^{-2})$. Since $\psi$ is decreasing with respect to $x$, we will consider when $\varepsilon_{0}(n)$ changes. More precisely, consider a minimal $k$ such that $\varepsilon_{0}(n)<\psi(n-k,\delta)\leqslant\varepsilon_{0}(n-k)$, which occurs to be the neat upper bound for $\varepsilon(n)$. However, since $\varepsilon_{0}(n)$ is the non-ascending step function, we realize that
$\varepsilon_{0}(n-k)\geqslant\left(2^{\left\lfloor-\log\left(1-\delta^{\frac{1}{n}}\right)\right\rfloor-1}-1\right)^{-1}\!\!\!\!=-\frac{2\ln(\delta)}{n}+\frac{3\ln(\delta)^{2}}{n^{2}}-\frac{13\ln(\delta)^{2}}{3n^{3}}+O(n^{-4})~{}.$
If we denote $\phi(n,\delta):=\left(2^{\left\lfloor-\log\left(1-\delta^{\frac{1}{n}}\right)\right\rfloor-1}-1\right)^{-1}$, then we can summarize our recent considerations in a short way by $\psi(n,\delta)\leqslant\varepsilon(n)<\phi(n,\delta)$. Thence, in the case where we fix the parameter $\delta=0.00033$, we obtain
$\frac{8.0164\ldots}{n}+\frac{32.13147\ldots}{n^{2}}+O(n^{-3})\leqslant\varepsilon(n)\leqslant\frac{16.0328\ldots}{n}+\frac{192.789\ldots}{n^{2}}+O(n^{-3})~{}.$
On the other hand, from Theorem 1, we know that when $\delta=0.00033$, then for Morris Counter (with $\varepsilon(n)$ defined by (10)) the quite similar relation holds:
$\varepsilon(n)\leqslant-\ln\left(1-\frac{16}{n}\right)=\frac{16}{n}+\frac{128}{n^{2}}+O(n^{-3})~{}.$
Therefore, Morris and MaxGeo Counters behave quite similarly under comparable conditions, and Figure 3 confirms this observation. Indeed, in Figure 3 we may see that the difference between the values of $\varepsilon(n)$ parameters for both counters decreases as $n$ increases.
Figure 3: Values of $\varepsilon(n)$ parameters for Morris and MaxGeo Counters compared with boundaries for $\varepsilon(n)$ for MaxGeo Counter: the lower one — $\psi(n,\delta)$ and the upper one — $\phi(n,\delta)$ ($n\leqslant 160$ and $\delta=0.00033$).
Realize that the previous conclusions remain true if $\delta(n)$ is not constant. This short observation enables us to obtain a more general result:
Fact 3
Let $\delta(n)=n^{-c}$ for some constant $c>0$. Then
$\varepsilon(n)\leqslant\phi(n,\delta(n))=\frac{2c\ln(n)}{n}+\frac{3c^{2}\ln(n)^{2}}{n^{2}}+O\left(\frac{c^{3}\ln(n)^{3}}{n^{3}}\right)$
and MaxGeo Counter is $(\phi(n,\delta(n)),\delta(n))$-DP for any $n\in\mathbb{N}$.
Notice that in this case, both sequences of parameters tend to 00, which may be used as an advantage in applications, especially when we expect that the total number of increment requests will be very large. However, we emphasize that this requires $\delta(n)$ to be negligible.
Method Laplace noise Morris counter $\mathrm{LogLog}^{(2)}$ counter
$\mathcal{L}(n/16)$ $M_{n}$ $(M_{n}[1],M_{n}[2])$
$(\varepsilon,\delta)$-DP $(16/n,0)$-DP $(\frac{16}{n-8},\delta)$-DP $(\sim\frac{32.066}{n},\delta)$-DP
Estimator $\hat{n}$ $n+\mathcal{L}(n/16)$ $22^{M_{n}}-$ 2 $\sim 0.89\cdot 2^{\frac{1}{2}\sum_{i=1}^{2}M_{n}[i]}$
$\mathrm{Var}(\hat{n})$ $\frac{n^{2}}{128}$ $\frac{n^{2}+n}{2}$ $\sim 0.845\cdot n^{2}$
Avg. memory $\log(n)+O(1)$ $\log(\log(n))+O(1)$ $2\log(\log(\frac{n}{2}))+O(1)$
Figure 4: A summary of data aggregation techniques. The standard one is based on the Laplace method, and the rest are based on probabilistic counters. Recall that $\delta=0.00033$.
In Figure 4, we may briefly see that probabilistic counters can be used for data aggregation in order to decrease the memory usage in exchange for a slight increase of $\delta$ parameter of differential privacy and wider confidence intervals (lower accuracy). In recent years Big Data related problems became very popular. Note that this kind of application makes major use of memory. When the server aggregates many different data, standard solutions may cause a serious problem with data storage, which can be encountered by using the idea based on a probabilistic counter instead.
Example 5
Imagine that 100 million people participate in a general health survey with $100$ yes/no sensitive questions.
For every question, we would like to estimate the number of people who answered yes, but we want to guarantee differential privacy property at a reasonable level. Realize that if the number of yes answers is very small for some questions (e.g., when the question is about a very rare disease), then the number of no answers may be counted instead. Obviously, this method gains privacy, but loses precision.
According to Figure 4, if we use the Laplace method, then we may need approximately $100\log(10^{8})=2657.54\ldots$ bits to store the counters. However, if we use Morris Counter instead, about $100\log(\log(10^{8}))=473.20\ldots$ bits are needed. Let us note that all terms $O(1)$ in the ”Average memory” row of Figure 4 are bounded by $1$. Hence their impact is negligible from a practical point of view.
One may also complain about the heavy use of Pseudo Random Number Generator that probabilistic counters make. However, this problem may be resolved by generating the number of increment requests, which have to be forgotten until the next update of the counter by utilizing appropriate geometric distributions (see, for instance, [46] for a similar approach applied to reservoir sampling algorithm). This way, the use of PRNG can be substantially reduced.
8 Proofs
8.1 Main Theorem for Morris Counter (Theorem 1)
The proof is complicated and very technical. In order to better understand it, we are going to provide a presentation of a plan and main ideas beneath the parts of the proof. Let $n_{k}:=2^{k}+1$ for $k\in\mathbb{N}$. We introduce $\mathcal{P}_{k}^{(c)}$ to be the probability of the Morris Counter $M_{n}$ being $k+c$ after $n_{k}$ increment requests, i.e. $p_{n_{k},k+c}$. The ”special” sequences $(\mathcal{P}_{k}^{(c)})_{k}$ play a crucial role in the proof.
Roadmap of the proof
We can divide the proof of Theorem 1 into five phases (the main results of the phases are given in brackets):
  1. 1.
    $\delta$ phase (Theorem 5),
  2. 2.
    relations between ”special” sequences $(\mathcal{P}_{k}^{(c)})_{k}$ with respect to $c$ (Claim 1),
  3. 3.
    dependencies between consecutive distributions (of $M_{n}$ and $M_{n+1}$) (Claim 2),
  4. 4.
    extrapolation of $\mathcal{P}_{k}^{(c)}\leq 2^{c+3}\mathcal{P}_{k}^{(c+1)}$ property to $N>n_{k}$ (Lemma 3),
  5. 5.
    $\varepsilon$ phase (Theorem 6).
During the first phase, we consider a concentration of Morris Counter in the vicinity of its mean value. More precisely, we show that Morris Counter after $n$ increment requests takes values in relatively small intervals $I_{n}$ with probability $1-\delta$ (note that then $M_{n}$ satisfies the condition (2) for $S_{n}=I_{n}$), where $I_{n}$ is defined as in (4) and $\delta$ is some small constant, which arises from the proof. Note that $I_{n}$ may be interpreted as confidence intervals at level $1-\delta$ (see e.g. [47]). This phase is divided into lemmas 1 and 2. The first one uniformly bounds the formula for probabilities given by the Theorem below, due to Flajolet:
Theorem 4 (Proposition 1 from [2])
The probability $p_{n,l}$ that the Morris Counter has value $l$ after $n$ increment requests is
$p_{n,l}=\sum_{j=0}^{l-1}\left(-1\right)^{j}2^{-j\left(j-1\right)/2}\left(1-2^{-\left(l-j\right)}\right)^{n}\prod_{i=1}^{j}\left(1-2^{-i}\right)^{-1}\prod_{i=1}^{l-1-j}\left(1-2^{-i}\right)^{-1}~{}.$
In Lemma 1, we sum up bounds on $p_{n,l}$ to obtain small upper bound for $\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)$. The same bounds cannot be efficiently utilized in the proof of Lemma 2. Instead, it couples $M_{n}$ with a process $X_{n}$, which increases during first $\left\lceil\log(n)\right\rceil+1$ steps, and then follow the same rule of update, so $M_{n}\leqslant X_{n}$ almost surely. Therefore $\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$, which is much easier to bound from definition. Let us note that such coupling cannot be used in the proof of Lemma 1. The first phase is summarized by Theorem 5, i.e., establishes $\delta=\delta_{1}+\delta_{2}$.
In the second phase we show that $(\mathcal{P}_{k}^{(4)})_{k}$ is descending for big enough $k$ and $(\mathcal{P}_{k}^{(5)})_{k}$ is ascending for big enough $k$ (Lemma 6). A change of monotonicity is the first obstacle in the proof of the main theorem. The main idea beneath the proof of this fact is to calculate the differences between consecutive elements of the considered ”special” sequences by representing them as the sums via application of Flajolet’s Theorem 4 and realizing that usually at most first ten terms of the sums are crucial (on the other hand, taking less than eight terms is rarely sufficient). This is the second issue, which makes the proof so complicated. Let us notice that Theorem 4 presents an explicit formula for $\mathbb{P}\left(M_{n}=l\right)$, which (as we may experience in A) is not convenient to analyze. However, it is simple enough to find the values numerically (also note that recursive Definition 3 provides those probabilities easily as well. However, this approach is inefficient in terms of memory and time for a big number of requests $n$). Therefore, by precise analysis, we can finally check some sums numerically and obtain the thesis of Lemma 6 for $k\geqslant 15$. However, numerically, one can extrapolate it to some smaller $k$ as well (remark that this proof does not work for small values of $k$, since not only ten terms of the aforementioned sums are important). Lemma 6 can be directly used to show that $\mathcal{P}_{k}^{(4)}\leqslant 2^{7}\mathcal{P}_{k}^{(5)}$ for $k\geqslant 7$ (Claim 1). Let us note that in fact, this result is not true for $k<7$.
The third phase is based on the foregoing intuition: If the ratios of consecutive probabilities of distribution of $M_{n}$ increase almost exponentially, then the ratios of distribution of $M_{n+1}$ increase similarly (but slightly slower). The proofs mainly use the definition of the Morris counter.
The fourth phase (i.e. Lemma 3) shows that if $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$, for $c\in[-k:4]$, then the same is true if we substitute $n_{k}$ with a bigger number (i.e., this property is increasing with respect to the $n$ parameter). In order to apply this result, we have to satisfy some starting conditions. We have numerically checked the appropriate condition for $k=7$ (presented later, in Table 1). Therefore, the second phase of the proof of Theorem 1 justifies the assumptions of Lemma 3 with respect to the $k$ parameter, as long as $k\geqslant 7$ and the third phase let us obtain the appropriate assumptions with respect to $c$. One can check that for $k<7$, an analogous assumption is not true.
The latter phase begins with the application of the result from the previous part. We obtain that $\varepsilon$ is at most $L(n)$ (provided in the formulation of Theorem 1), for $k\geqslant 7$. The last piece of this puzzle is justified by a numerical evaluation for $k<7$ (presented later, in Figure 5), which ends the $\varepsilon$ phase, and so the whole proof.
Proof
In the following, we present the proof of the main contribution. Nevertheless, some technical lemmas are given in the A.
$\delta$ phase
Let us begin with a reminder. First of all, $M_{n}$ is ranged in $\mathbb{N}_{0}$ and, moreover, $I_{n}\subseteq\left[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4\right]$. We provide few facts about the concentration of the distribution of the random variable $M_{n}$, or more precisely about the probability that $M_{n}$ will be outside the interval $I_{n}$.
Lemma 1
Let $M_{n}$ be the state of the Morris counter after $n$ increment requests. Then
$\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)\leqslant 0.000006515315\ldots~{}.$
An increasing sequence $\prod\limits_{i=1}^{k}\left(1-2^{-i}\right)^{-1}$ that arose in Theorem 4 will be indicated by $r_{k}$ (with $r_{0}=1$) and we denote its limit $\prod\limits_{i=1}^{\infty}\left(1-2^{-i}\right)^{-1}=3.46274\ldots$ by $R$.
Let us mention that with the notions $q_{k}$ and $Q$ from [2], $r_{k}=\frac{1}{q_{k}}$ and $R=\frac{1}{Q}$.
Proof 1
At first, we want to bound a lower tail of the distribution $\delta_{1}$. Here we would like to find a sufficient upper limit for the above probability. Assume that $l\leqslant\left\lceil\log(n)\right\rceil-5$. Let us realize that $r_{k}\leqslant R$ and that $y=-x+1$ is a tangent line to the plot of $y=-x(x-1)$ in the point $(x,y)=(1,0)$. Therefore:
[…]
$\leqslant R^{2}\frac{2}{\sqrt{2}-1}\exp(-n2^{-l})=R^{2}(2\sqrt{2}+2)\exp(-n2^{-l}).$
Above formula will help us to limit the left tail od the distribution of $M_{n}$:
$\delta_{1}=\sum_{l=1}^{\left\lceil\log(n)\right\rceil-5}\mathbb{P}\left(M_{n}=l\right)\leqslant R^{2}(2\sqrt{2}+2)\sum_{l=1}^{\left\lceil\log(n)\right\rceil-5}\exp(-n2^{-l})\\ \leqslant R^{2}(2\sqrt{2}+2)\sum_{k=4}^{\infty}\exp(-2^{k})\leqslant R^{2}(2\sqrt{2}+2)\sum_{k=1}^{\infty}\exp(-16k)\\ =R^{2}(2\sqrt{2}+2)\frac{\exp(-16)}{1-\exp(-16)}=0.000006515315\ldots~{}.$
$\qed$
Remark
The bound for $p_{n,l}$ obtained above, is useless when $l\geqslant\log(n)-2$, so it cannot be used in the next lemma for an symmetrical upper tail.
Lemma 2
Let $M_{n}$ be the state of the Morris Counter after $n$ increment requests. Then
$\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant 0.000325521\ldots~{}.$
Proof 2
Consider a process $X=(X_{k},k\in\![0:n])$. Let $X$ initially follow the incrementation rule $\mathbb{P}\left(X_{k}=k+1\right)=1$ for $k\in[0:\lceil\log n\rceil+1]$. Afterwards, let this Markov chain imitate the transition rule of Morris Counter, that is
$\mathbb{P}\left(X_{k+1}=m+1|X_{k}=m\right)=\frac{1}{2^{m}}=1-\mathbb{P}\left(X_{k+1}=m|X_{k}=m\right)$
for $k\geqslant\left\lceil\log(n)\right\rceil+1$. Naturally, for $k\leqslant\left\lceil\log(n)\right\rceil+1$, we have $X_{k}\geqslant M_{k}$, so we may couple realizations of these two processes in such a way that whenever $X$ is incremented, then so is $M$ and if $M$ does not change, then $X$ does not increase as well (note that $X$ has at most the same probability of a positive increment as $M$ at any point of time).
To abbreviate the expressions let us denote $m=n-\left\lceil\log(n)\right\rceil-1$ and
$\mu_{\iota}=\mathbb{P}\left(X_{k+1}=\left\lceil\log(n)\right\rceil+\iota+1|X_{k}=\left\lceil\log(n)\right\rceil+\iota\right)=\frac{1}{2^{\left\lceil\log(n)\right\rceil+\iota}}=1-\nu_{\iota}~{},$
for any $\iota\in\mathbb{Z}$. Moreover, let us define a three-dimensional discrete simplex:
$S_{k}^{(3)}=\{\bar{l}=(l_{1},l_{2},l_{3})\in\mathbb{N}_{0}^{3}:l_{1}+l_{2}+l_{3}\leqslant k\}~{}.$
$\delta_{2}\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$ $=\sum_{\bar{l}\in S_{m-3}^{(3)}}\nu_{2}^{l_{1}}\mu_{2}\nu_{3}^{l_{2}}\mu_{3}\nu_{4}^{l_{3}}\mu_{4}\leqslant\sum_{\bar{l}\in S_{m-3}^{(3)}}\dfrac{1}{2^{3\left\lceil\log(n)\right\rceil+9}}$
$=\sum_{k=0}^{m-3}\binom{k+3}{2}\dfrac{1}{2^{3\left\lceil\log(n)\right\rceil+9}}\leqslant\frac{1}{2^{10}n^{3}}\sum_{k=3}^{m}k^{2}-k~{}.$
Realize that $\sum\limits_{k=3}^{m}k=(m-2)(m+3)/2$ and $\sum\limits_{k=3}^{m}k^{2}=(m-2)(2m^{2}+7m+15)/6$, so
$\delta_{2}$ $\leqslant\frac{1}{2^{10}n^{3}}\frac{1}{6}(m-2)(2m^{2}+4m+6)=\frac{1}{3\cdot 2^{10}n^{3}}(m^{3}-m-6)$
$\leqslant\frac{m^{3}}{3\cdot 2^{10}n^{3}}\leqslant\frac{1}{3\cdot 2^{10}}=0.000325521\ldots~{}.$
Note that when $m<3$ (that is, when $n<7$), then the above sums are empty, but on the other hand $\lceil\log(n)\rceil+5>n+1$, so the inequality is trivially true.$\qed$
Theorem 5
The state of the Morris Counter after $n$ increment requests is not in the set
$I_{n}=[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4]\cap[n+1]~{}$
with probability $\delta<0.00033$.
Proof 3
Realize that $\mathbb{P}\left(M_{n}\in[1:n+1]\right)=1$. This observation, together with lemmas 1 and 2 yield
$\delta:=\mathbb{P}\left(M_{n}\notin I_{n}\right)=\delta_{1}+\delta_{2}<0.00033~{}.$
$\qed$
The second phase
In this part of the investigation, we try to establish $\varepsilon(n)$ parameter of DP of $M_{n}$. In fact it remains to examine the property (1) in the interval $I_{n}$, as Theorem 5 entails (2) for $S_{n}=I_{n}$. Therefore, we are interested in finding the upper bound for maximal privacy loss for any $n\in\mathbb{N}$ and $k\in I_{n}$, namely:
$\varepsilon(n)=\max\left\{\left|\ln\left(\frac{p_{n\pm 1,k}}{p_{n,k}}\right)\right|:k\in I_{n}\right\}~{}.$ (10)
Actually, we may consider ${}^{\prime}+^{\prime}$ sign instead of ${}^{\prime}\pm^{\prime}$ in (10), because $|\ln(x)|=|\ln({1}/{x})|$. However, when $I_{n}\neq I_{n\pm 1}$, we have to behave carefully, so in particular, an additional check of privacy loss with ${}^{\prime}-^{\prime}$ sign is needed when $n$ is of a form $12^{l}+$ 1 for some $l\in\mathbb{N}$.
Claim 1
For $k\geqslant 7$, we have $p_{2^{k}+1,k+4}\leqslant 2^{7}p_{2^{k}+1,k+5}$.
The above claim is the result of a simple application of Lemma 6 from A.
The third phase
Claim 2
If for any given $n$, there exists an ascending and positive sequence $(\alpha_{i})_{i=1}^{n}$ such that $(\forall\;i\in[1:n])\;p_{n,i}=2^{i}\alpha_{i}p_{n,i+1}$, then there also exists an ascending and positive sequence $(\alpha_{i}^{\prime})_{i=1}^{n+1}$ such that
$(\forall\;i\in[1:n+1])\;(p_{n+1,i}=2^{i}\alpha_{i}^{\prime}p_{n+1,i+1})\wedge(\forall\;i\in[1:n])\;(\alpha_{i}^{\prime}<\alpha_{i})~{}.$
This claim emerges from lemmas 7 and 8 from A.
The fourth phase
We use Claim 2 to guarantee starting conditions for the next Lemma 3. However, in order to apply Lemma 3, we will use Claim 1 as well, which assumes that $n\geqslant 2^{7}+1$. Hence we would like to gather some information about the distribution of $M_{2^{7}+1}$. More precisely, we are interested in the behaviour of $\theta_{i}={p_{129,i}}/{p_{129,i+1}}$ for $i\leqslant 11$, which we present in the Table 1.
i $\theta_{i}$ $2^{i-4}$ $2^{4-i}\theta_{i}$
1 $9.6205\ldots\cdot 10^{-24}$ 0.125 $7.6964\ldots\cdot 10^{-23}$
2 $1.73351\ldots\cdot 10^{-9}$ 0.25 $6.93402\ldots\cdot 10^{-9}$
3 $0.000119359\ldots$ 0.5 $0.000238718\ldots$
4 $0.0140238\ldots$ 1 $0.0140238\ldots$
5 $0.158163\ldots$ 2 $0.0790814\ldots$
6 $0.771817\ldots$ 4 $0.192954\ldots$
7 $2.67702\ldots$ 8 $0.334628\ldots$
8 $7.83367\ldots$ 16 $0.489604\ldots$
9 $20.8095\ldots$ 32 $0.650297\ldots$
10 $52.0472\ldots$ 64 $0.813238\ldots$
11 $125.065\ldots$ 128 $0.977073\ldots$
Table 1: Ratios of adjacent probabilities of the distribution of $M_{2^{7}+1}$, compared with the exponential function of base $2$.
We briefly see a superexponential trend of proportions $\theta_{i}$, so a possibility to use Claim 2 for $n\geqslant 2^{7}+1$ is justified. It might seem that the choice of $n$ is arbitrary, but it occurs that the distribution of $M_{2^{6}+1}$ does not have satisfy the necessary assumptions for the privacy loss although $M_{2^{6}+1}$ still can fulfill the property of $(\varepsilon(n),\delta)$-DP with parameters given in Theorem 1.
Lemma 3
Let $k\in\mathbb{N}\setminus\{0\}$ and $n_{k}=2^{k}+1$. If $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$ for every $c$ in the interval $[-k:4]$, then
$(\forall\;N\geqslant n_{k})(\forall\;c\in[-k:4])\;p_{N,k+c}<2^{c+3}p_{N,k+c+1}~{}.$
Proof 4
Realize that for $c=-k$, the required inequality is trivial. Therefore we can safely consider only $c\in[-k+1:4]$. We would like to prove it inductively with respect to $c$ and $N$. Assume that for some $N\geqslant n_{k}$ and any $d\in\{c-1,c\}$ we have $p_{N,k+d}<2^{d+3}p_{N,k+d+1}$. Then also
[…]
$\leqslant 2^{3+c}p_{N,k+c+1}(1-2^{-k-c})+2^{3+(c-1)}p_{N,k+c}2^{-k-c+1}$
$<2^{3+c}(p_{N,k+c+1}(1-2^{-k-c-1})+2^{-k-c}p_{N,k+c})=2^{3+c}p_{N+1,k+c+1}~{}.$
If we start with $c=-k+1$, then the above let us prove inductively the appropriate condition for all $N\geqslant n_{k}$. Further the thesis is followed by the induction with respect to $c$$\qed$
$\varepsilon$ phase
Claims 1 and 2, together with Table 1 enable us to apply Lemma 3 for $n=2^{k}+1$ for any $k\geqslant 7$.
Theorem 6
Let $n>2^{7}=128$ and $k\in I_{n}$. Then
$1-\frac{16}{n}\leqslant\frac{p_{n\pm 1,k}}{p_{n,k}}\leqslant 1+\frac{16}{n}~{}.$
Proof 5
According to the previous discussion about the formula (10), we examine:
[…]
Let us denote $l=\left\lceil\log(n)\right\rceil$ and $c=k-l\in[-4:4]$. Then Lemma 3 bears $p_{n,k-1}\leqslant~{}2^{c+3}p_{n,k}$, so
$\frac{p_{n+1,k}}{p_{n,k}}\leqslant 1+2^{-l-c}(-1+2^{c+4})<1+2^{-l+4}=1+\frac{16}{2^{\lceil\log(n)\rceil}}<1+\frac{16}{n}~{}.$
Realize that if $n=2^{l-1}+1$ for some $l\in\mathbb{N}$, then a little adjustment is necessary. Indeed, let now $c-1=k-l\in[-4:4]$, and once again, Lemma 3 provides $p_{n-1,k-1}<~{}2^{c+2}p_{n-1,k}$. However, it still holds that:
$\frac{p_{n,k}}{p_{n-1,k}}=1+2^{-k}\left(-1+2\frac{p_{n-1,k-1}}{p_{n-1,k}}\right)\leqslant 1+2^{-l-c+1}\left(-1+2^{c+3}\right)<1+\frac{16}{n}~{}.$
On the other hand, we obviously have inequalities $p_{n+1,k}>\left(1-2^{-l-c}\right)p_{n,k}$ and $p_{n,k}>\left(1-2^{-l-c}\right)p_{n-1,k}$ for any $c\in[-4:4]$, so both of these fractions exceed $1-{16}/{n}$$\qed$
Theorem 6 only provides $\varepsilon(n)\leqslant-\ln\left(1-{16}/{n}\right)$ for $n>128$ (compare with (10)). However, in the Figure 5 we may briefly see that the above inequality is true for smaller number of requests $n$ as well.
Having all the technical lemmas, we are now ready to prove Theorem 1.
Proof 6
(of Theorem 1) Suppose that $S_{n}=I_{n}$ in Fact 1. Then from theorems 5 and 6 we can easily see that $\mathbb{P}\left(M_{n}\notin S_{n}\right)<0.00033$ and
$(\forall\;m\in\{n-1,n+1\})\;(\forall\;l\in S_{n})\;\mathbb{P}\left(M_{n}=l\right)\leqslant\left(1-\frac{16}{n}\right)^{-1}\cdot\mathbb{P}\left(M_{m}=l\right)~{},$
hence from Fact 1 we obtain the main result.$\qed$
Figure 5: Exact values of $\varepsilon(n)$ parameter for $n\leqslant 160$ compared with plots of sequences $-\ln(1-{16}/{n})$ and $-\ln(1-{8}/{n})$.
In the Figure 5 one can see that values of $\varepsilon(n)$ are strictly between sequences $-\ln(1-~{}{8}/{n})$ and $-\ln(1-~{}{16}/{n})$ for $n\in[17:160]$. We can also observe that $\varepsilon(n)~{}\approx~{}2^{4-\lceil\log(n)\rceil}$ in this interval. Note that $\lceil\log(n)\rceil\leqslant 4$ for $n\leqslant 16$, so $\lceil\log(n)\rceil-4<1$, but $M$ is always positive. This can justify the chaotic behavior of the process for $n\leqslant 16$. Nevertheless, Figure 5 affirms the quality of $\varepsilon(n)$ parameter established in Theorem 1. Moreover, let us mention, that for the strategy used in the proof, we cannot pick smaller $\varepsilon$ of the same form:
Observation 2
Assume that for $I_{n}$, defined in (4), $\delta=\Pr{M_{n}\notin I_{n}}$ and $\varepsilon(n)$, defined via (10), is of the form $|-\ln(1-\frac{c}{n})|$. Then $c=16$ is optimal constant in Theorem 1 (it cannot be improved). Indeed, see that the bound is reached for $n=32$:
$\dfrac{p_{33,1}}{p_{32,1}}=\frac{1}{2}=1-\frac{16}{32}~{}.$
As may be seen in Figure 5, minimal $c$ for $\varepsilon(n)$, where $n$ is of the form $2^{k}$, for $k>5$, is very close to $16$ as well.
8.2 General result on Morris’ Counter privacy (Theorem 2)
Remark that $\lceil c\log(\ln(n))\rceil\geqslant 1$ can be always guaranteed, when the $n$ is big enough.
Proof 7
For our convenience, let us denote $\rho:=\lceil c\log(\ln(n))\rceil$. We assume that $\rho\geqslant 1$. First we show that $\delta_{1}^{\ast}:=\mathbb{P}\left(M_{n}\leqslant\lceil\log(n)\rceil-\rho-1\right)=O\left(n^{-\left(\ln(n)\right)^{c-1}}\right)$. The proof is analogous to the one of Lemma 1 (we omit the similar parts). Indeed,
$\delta_{1}^{\ast}=\sum_{l=1}^{\left\lceil\log(n)\right\rceil-\rho-1}\mathbb{P}\left(M_{n}=l\right)~{}\leqslant R^{2}(2\sqrt{2}+2)\sum_{k=\rho}^{\infty}\exp(-2^{k})\\ \leqslant R^{2}(2\sqrt{2}+2)\sum_{k=1}^{\infty}\exp(-2^{\rho}k)\\ \leqslant R^{2}(2\sqrt{2}+2)\frac{\exp(-(\ln(n))^{c})}{1-\exp(-(\ln(n))^{c})}=O\left(n^{-\left(\ln(n)\right)^{c-1}}\right)~{}.$
Now, we are going to prove $\delta_{2}^{\ast}:=\mathbb{P}\left(M_{n}\geqslant\lceil\lg(n)\rceil+\rho+1\right)=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)$. We use similar notation and technique as in proof of Lemma 2, but this time we utilize $(\rho-1)$-dimensional discrete simplex:
$S_{k}^{(\rho-1)}=\left\{\bar{l}=(l_{1},l_{2},\ldots,l_{\rho-1})\in\mathbb{N}_{0}^{\rho-1}:\sum_{i=1}^{\rho-1}l_{i}\leqslant k\right\}~{}.$
We couple $(M_{n})_{n}$ with the same process $(X_{n})_{n}$ as in Lemma 2. Roughly speaking, $X_{0}=1$ and $X_{n}$ almost always increments by $1$ for until $n=\lceil\log(n)\rceil+1$ and further it follows the same rule of incrementation as Morris counter. Then
$\delta_{2}^{\ast}$ $\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+\rho+1\right)=\sum_{\bar{l}\in S_{m-\rho+1}^{(\rho-1)}}\prod_{i=1}^{\rho-1}\nu_{i+1}^{l_{i}}\mu_{i}\leqslant\sum_{\bar{l}\in S_{m-\rho+1}^{(\rho-1)}}\prod_{i=1}^{\rho-1}\mu_{i}$
$=\sum_{\bar{l}\in S_{m-\rho+1}^{(\rho-1)}}2^{-[(\rho-1)\left\lceil\log(n)\right\rceil+\sum_{i=2}^{\rho}i]}$
$=\sum_{k=0}^{m-\rho+1}\binom{k+\rho-1}{\rho-2}2^{-[(\rho-1)\left\lceil\log(n)\right\rceil+\frac{(\rho+2)(\rho-1)}{2}]}$
$\leqslant m^{\rho-2}n^{1-\rho}2^{-\frac{\rho^{2}+\rho-2}{2}}\leqslant n^{-1}2^{-\rho+1}=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)~{}.$
Therefore $\mathbb{P}\left(M_{n}\notin J_{n}(c)\right)=\delta_{1}^{\ast}+\delta_{2}^{\ast}=O\left(n^{-\left(\ln(n)\right)^{c-1}}+n^{-1}\left(\ln(n)\right)^{-c}\right)$
Further we would like to consider fractions $\frac{p_{n+1,k}}{p_{n,k}}$ for $k\in J_{n}(c)$ as in the proof of Theorem 6. Indeed
$1-2^{-k}\leqslant\frac{p_{n+1,k}}{p_{n,k}}=1-2^{-k}+2^{-k+1}\frac{p_{n,k-1}}{p_{n,k}}~{}.$
We are going to use another formula from [2]. For any $n\in\mathbb{N}$ and $k\in[1:n+1]$,
$p_{n,k}=2^{-\frac{k(k-1)}{2}}\sum_{\bar{l}\in S_{n-k+1}^{(k)}}\prod_{i=1}^{k}\left(1-2^{-i}\right)^{l_{i}}~{}.$
Let us denote the above sum by $\varsigma_{k}(n-k+1)$. Let us note that
$2^{-k+1}\frac{p_{n,k-1}}{p_{n,k}}=2^{-k+1}\frac{2^{-\frac{(k-2)(k-1)}{2}}\varsigma_{k-1}(n-k+2)}{2^{-\frac{k(k-1)}{2}}\varsigma_{k}(n-k+1)}=\frac{\varsigma_{k-1}(n-k+2)}{\varsigma_{k}(n-k+1)}~{}.$
Realize that $\varsigma_{k-1}(n-k+2)\leqslant\varsigma_{k-1}(n-k+1)\sum_{i=1}^{k-1}\left(1-2^{-i}\right)=\varsigma_{k-1}(n-k+1)\left(k-2+2^{k-1}\right)$. This follows from the fact that each summand of $\varsigma_{k-1}(n-k+2)$ can be obtained from some summands of $\varsigma_{k-1}(n-k+1)$ by multiplication by one of the terms $\left(1-2^{-i}\right)$. Moreover, note that $\varsigma_{k}(n-k+1)$ has $\binom{n-k+1+(k-1)}{k-1}=\binom{n}{k-1}$ summands and similarly, $\varsigma_{k-1}(n-k+1)$ has $\binom{n-1}{k-2}$ summands. One can briefly see that a function $f(i)=\left(1-2^{-i}\right)$ is increasing, hence
$\varsigma_{k}(n-k+1)$ $\geqslant\sum_{\bar{l}\in S_{n-k+1}^{(k)}}\left(1-2^{-k+1}\right)^{l_{k-1}+l_{k}}\prod_{i=1}^{k-2}\left(1-2^{-i}\right)^{l_{i}}$
$=\sum_{\bar{l}\in S_{n-k+1}^{(k-1)}}(l_{k-1}+1)\prod_{i=1}^{k-1}\left(1-2^{-i}\right)^{l_{i}}~{}.$
Due to the monotonicity of $f$, one can use cascading substitutions: some of $f(k)$ by $f(k-1)$, then some of $f(k-1)$ by $f(k-2)$ etc., in order to balance the numbers of all the occurring summands, what provides:
$\varsigma_{k}(n-k+1)\geqslant\sum_{\bar{l}\in S_{n-k+1}^{(k-1)}}\frac{\binom{n}{k-1}}{\binom{n-1}{k-2}}\prod_{i=1}^{k-1}\left(1-2^{-i}\right)^{l_{i}}=\frac{n}{k-1}\varsigma_{k-1}(n-k+1)~{}.$
Therefore $\varsigma_{k-1}(n-k+2)\leqslant\frac{\left(k-2+2^{k-1}\right)(k-1)}{n}\varsigma_{k}(n-k+1)$ and finally we obtain
$\frac{p_{n+1,k}}{p_{n,k}}\leqslant 1-2^{-k}+\frac{\left(k-2+2^{k-1}\right)(k-1)}{n}<1+\frac{(k-1)^{2}}{n}~{}.$
When $k\in J_{n}(c)$, then
$\exp\left(-O\left(\frac{\log(n)}{n}\right)\right)=1-2^{-\lceil\log(n)\rceil+\lceil\log\ln(n)\rceil}\\ <\frac{p_{n+1,k}}{p_{n,k}}<\\ 1+\frac{(\lceil\log(n)\rceil+\log\ln(n))^{2}}{n}=\exp\left(O\left(\frac{(\log(n))^{2}}{n}\right)\right)~{}.$
This shows that $\varepsilon(n)=O\left(\frac{(\log(n))^{2}}{n}\right)$$\qed$
8.3 Main result for MaxGeo Counter (Theorem 3)
Proof 8
We have $n$ increment requests, which influence the value of MaxGeo Counter $M$. Thence the result of the mechanism is $X=\max(X_{1},\ldots,X_{n})$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. First, we observe that if $n=m$, the counter trivially satisfies differential privacy, as the probability distribution of $X$ does not change. From now on, we assume that $|n-m|=1$. See that
$\mathbb{P}\left(X\leqslant l\right)=\prod_{i=1}^{n}\mathbb{P}\left(X_{i}\leqslant l\right)=\left(\mathbb{P}\left(X_{1}\leqslant l\right)\right)^{n}=\left(1-\frac{1}{2^{l}}\right)^{n}=\left(\frac{2^{l}-1}{2^{l}}\right)^{n}~{}.$
Furthermore
$\mathbb{P}\left(\max(X_{1},\ldots,X_{n})=l\right)=\mathbb{P}\left(X=l\right)=\mathbb{P}\left(X\leqslant l\right)-\mathbb{P}\left(X\leqslant(l-1)\right)$
$=\left(\frac{2^{l}-1}{2^{l}}\right)^{n}-\left(\frac{2^{l-1}-1}{2^{l-1}}\right)^{n}=\frac{\left(2^{l}-1\right)^{n}-\left(2^{l}-2\right)^{n}}{2^{l\cdot n}}~{}.$
Now we need to calculate the following expression
$\frac{\mathbb{P}\left(\max(X_{1},...X_{n})=l\right)}{\mathbb{P}\left(\max(X_{1},...X_{n},X_{n+1})=l\right)}$ $=\dfrac{\dfrac{\left(2^{l}-1\right)^{n}-\left(2^{l}-2\right)^{n}}{2^{l\cdot n}}}{\dfrac{\left(2^{l}-1\right)^{n+1}-\left(2^{l}-2\right)^{n+1}}{2^{l\cdot(n+1)}}}$
$=\frac{\left(\left(2^{l}-1\right)^{n}-(2^{l}-2)^{n}\right)\cdot 2^{l}}{\left(2^{l}-1\right)^{n+1}-(2^{l}-2)^{n+1}}$
$=\frac{2^{l}}{2^{l}-1}\cdot\frac{\left(\left(2^{l}-1\right)^{n}-(2^{l}-2)^{n}\right)}{\left(\left(2^{l}-1\right)^{n}-\frac{(2^{l}-2)^{n+1}}{2^{l}-1}\right)}$
$\leqslant\frac{2^{l}}{2^{l}-1}\cdot\frac{\left(\left(2^{l}-1\right)^{n}-(2^{l}-2)^{n}\right)}{\left(\left(2^{l}-1\right)^{n}-\frac{(2^{l}-2)^{n+1}}{2^{l}-2}\right)}$
$=\frac{2^{l}}{2^{l}-1}=1+\frac{1}{2^{l}-1}~{}.$
For fixed $\varepsilon$ we need to satisfy the following inequality
$\left|\ln\left(\frac{P(\max(X_{1},...X_{n})=l)}{P(\max(X_{1},...X_{n},X_{n+1})=l)}\right)\right|\leqslant\varepsilon~{},$
which gives
$\ln\left(1+\frac{1}{2^{l}-1}\right)\leqslant\varepsilon~{}.$ (11)
We can see from (11) that the greater $l$ is, the smaller $\varepsilon$ can be. Moreover, inequality (11) is true for $l\geqslant l_{\varepsilon}$. Therefore, we must ensure $P(X\leqslant l_{\varepsilon})\leqslant\delta$. See that
$\mathbb{P}\left(X\leqslant l_{\varepsilon}\right)=\left(1-2^{-l_{\varepsilon}}\right)^{n}~{}.$
It is easy to see that the above decreases when $n$ increases. Then
$\left(1-2^{-l_{\varepsilon}}\right)^{n}\leqslant\delta\iff n\geqslant\frac{\ln(\delta)}{\ln(1-2^{-l_{\varepsilon}})}\approx-\frac{\ln(\delta)}{\varepsilon},$
where the approximation is a result of substitution of $l_{\varepsilon}$ without ceiling. $\qed$
9 Conclusions and Future Work
In this paper, we have investigated probabilistic counters from the privacy-protection perspective. We have shown that Morris Counter and MaxGeo Counter inherently guarantee differential privacy from the mechanism itself, provided that there is at least a small, fixed number of increment requests. Otherwise, the counter has too low value, and intuitively, the result is not randomized enough. We have also shown that the constant in our Morris Counter result cannot be improved further.
We have shown how to perform data aggregation, namely a distributed survey, in a privacy-preserving manner using probabilistic counters. We clarified that this type of solution is especially efficient when one cares about memory resources, like in many Big Data related problems. Note that the security model in this paper was somewhat optimistic. Unfortunately, in such a setting, there is little incentive to use them other than when we already have them deployed and working as aggregators due to e.g., memory-efficiency requirements. This would, however, change tremendously if we would weakened these assumptions. This seems a very promising way to continue our research from this paper. Namely, we focused on privacy and can still not weaken the security assumptions and allow the Adversary to extract information from channels between users and the aggregator. That would put us in the so-called, Local Model, where each user is responsible for the data randomization. However, such an approach requires us to be able to perform probabilistic counter in an oblivious manner which, to the best of our knowledge, was not explored before.
In Subsection 4.1, we have mentioned the generalization of Morris Counter (for bases $a>1$). Analysis of privacy properties of such variants of Morris Counters and various probabilistic counters presented, for example, in [28], [48] may also be promising direction of further research.
In this paper, we focused on the standard definition of differential privacy. However, there is also an issue of preservation of differential privacy for requests given by a group of $k$ users or one individual sending up to $k$ dependent requests over time. This can be described in a language of so-called $(\varepsilon,\delta)-k$-DP (see "group privacy scenario" in [5]) . A group of people may tend to behave in the same manner, so they may send $k$ requests in a row. Especially this ”group” may be represented by a single person colluding with the Adversary. It is worth mentioning that this type of generalization creates an opportunity to modify probabilistic counters so that each incrementation request executes the update request multiple times to reduce the variance of the rescaled estimator. Intuitively, this extension should be especially efficient in preserving standard differential privacy property when $\varepsilon(n)=\frac{c}{n}+o(n^{-1})$ (as a parameter of standard differential privacy), because both $c$ and $n$ should scale with $k$ linearly. Hence the next challenging problem is showing that Morris and MaxGeo Counters satisfy $k$-DP property with similar privacy parameters.
Morris Counter and MaxGeo Counter are considered the most popular probabilistic counters. However, the results of this paper shed new light on the properties of the probabilistic counter in general. There is a possibility to provide analogous differential privacy properties for other probabilistic counters. Moreover, this paper enables the provision of differentially private algorithms for other applications, especially those based on Morris or MaxGeo Counter. For instance, in Section 4 we mentioned PCSA and HyperLogLog Counters together with their variances, which can be manually adjusted to applications. A proper choice of $m$ parameter implies an exchange of the memory usage for the accuracy of the estimation. We have mentioned that these counters’ differential privacy parameters can be obtained via Observation 1. However, such a direct result may not be satisfying. Hence a more precise calculation is needed. For instance, Observation 1 may be utilized again with some concentration inequalities.
Acknowledgement
This work was supported by Polish National Science Center grant number UMO-2018/29/B/ST6/02969.
Appendix A Technical Lemmas and Proofs Related to Differential Privacy of Morris Counter
For the sake of completeness, we present here proofs of all technical lemmas that are not directly connected to Theorem 1. Some of computations are supported by Wolfram Mathematica ver.11.3 ([49]). Whenever we obtain a result in this manner, we indicate it by =$\stackrel{{\scriptstyle W}}{{=}}$ sign. Usually results are precise, however in some cases, final forms are attained numerically. We often struggle with expressions of a pattern $1-{1}/{y}$, so let us denote this function as $a(y)$ to abbreviate formulas.
Next two lemmas will be useful in a proof of Lemma 6.
Lemma 4
Let $c>{1}/{x}$. Then
$a(2cx)^{2y}\geqslant a(cx)^{y-1}\left(a(cx)+\frac{y}{4c^{2}x^{2}}\right)$
and
$a(cx)^{y}\geqslant a(2cx)^{2y-2}\left(a(2cx)^{2}-\frac{y}{4c^{2}x^{2}}\right)~{}.$
Proof 9
$\frac{a(2cx)^{2y}-a(cx)^{y}}{a(2cx)^{2}-a(cx)}=\sum_{i=0}^{y-1}a(2cx)^{2i}a(cx)^{y-i-1}~{}.$
Realize that the above denominator is $\left(1-\frac{1}{cx}+\frac{1}{4c^{2}x^{2}}-1+\frac{1}{cx}\right)=\frac{1}{4c^{2}x^{2}}$. Hence, we obtained two inequalities: $a(2cx)^{2y}-a(cx)^{y}\geqslant\dfrac{y}{4c^{2}x^{2}}a(cx)^{y-1}$
and $a(2cx)^{2y}-~{}a(cx)^{y}\leqslant~{}\dfrac{y}{4c^{2}x^{2}}~{}a(2cx)^{2(y-1)}$, which imply the thesis of this Lemma. $\qed$
Lemma 5
Let $s\leqslant\log({x}/{4})$. Then $a(2^{-s}x)^{2x+1}<\exp(-2^{s+1})$ and
$a(2^{-s}x)^{x-1}>\exp(-2^{s})\left(1-\frac{2^{2s-1}-2^{s}}{x}-\frac{2^{2s-7}+2^{4s-3}}{x^{2}}\right)~{}.$
Proof 10
Let $f_{1}(x;s):=a(2^{-s}x)^{2x+1}$. For any $s\leqslant\log(x/4)$, we have $f_{1}(x;s)=\left(1-\frac{2^{2+1}}{2x}\right)^{2x=1}=\exp(-2^{s+1})\left(1-O\left(x^{-1}\right)\right).$ Realize a fact, that $z\ln(z)\geqslant z-1$, for $0<z\leqslant 1$. Hence
$\left(1-\frac{2^{s}}{x}\right)^{-2x}\dfrac{\partial f_{1}(x;s)}{\partial x}=\frac{2^{s}(2x+1)}{x^{2}}+2\left(1-\frac{2^{s}}{x}\right)\ln\left(1-\frac{2^{s}}{x}\right)>\frac{2^{s}}{x^{2}}>0$
and in a consequence $a(2^{-s}x)^{2x+1}<\exp(-2^{s+1})$ for any reasonable $s$. Moreover, let us introduce
$D(x;s):=1-\frac{2^{2s-1}-2^{s}}{x}-\frac{2^{2s-7}+2^{4s-3}}{x^{2}}~{}.$
It is defined in such the way that $a(2^{-s}x)^{x-1}=D(x;s)+O(x^{-2})$. Therefore we can attain:
$f_{2}(x;s):=\dfrac{a(2^{-s}x)^{x-1}}{D(x;s)}=\exp(-2^{s})\left(1+O\left(x^{-2}\right)\right)~{}.$
Then, in a similar way
$D(x;s)^{2}\left(1-\frac{2^{s}}{x}\right)^{-x+1}\dfrac{\partial f_{2}(x;s)}{\partial x}=$
$D(x;s)\left(\frac{2^{s}(x-1)}{x^{2}\left(1-\frac{2^{s}}{x}\right)}+\ln\left(1-\frac{2^{s}}{x}\right)\right)-\left(\frac{2^{2s-6}+2^{4s-2}}{x^{3}}+\frac{2^{2s-1}-2^{s}}{x^{2}}\right)$
$<\left(\frac{2^{2s-1}-2^{s}}{x^{2}}+\frac{2^{3s}-2^{2s}}{x^{3}\left(1-\frac{2^{s}}{x}\right)}-\frac{2^{3s}}{3x^{3}}\right)-\left(\frac{2^{2s-6}+2^{4s-2}}{x^{3}}+\frac{2^{2s-1}-2^{s}}{x^{2}}\right)$
$=\frac{2^{3s}-2^{2s}}{x^{3}\left(1-\frac{2^{s}}{x}\right)}-\frac{2^{3s}}{3x^{3}}-\frac{2^{2s-6}+2^{4s-2}}{x^{3}}~{}.$
Let $d:=1-{2^{s}}/{x}$ and realize that $d\in[{3}/{4},1)$ and $02^{s}-1+d\left(-{2^{s}}/{3}-2^{-6}-2^{2s-2}\right)>$ 0. Indeed, if we put $z=2^{s}$, then we attain a quadratic inequality in $z$ variable, with determinant $\Delta=1-\frac{5d}{3}+\frac{55d^{2}}{576}$, that is negative for $d\in[{3}/{4},1)$.
Hence $\dfrac{\partial f_{2}(x;s)}{\partial x}<0$ and consequently
$a(2^{-s}x)^{x-1}>\exp(-2^{s})\left(1-\frac{2^{2s-1}-2^{s}}{x}-\frac{2^{2s-7}+2^{4s-3}}{x^{2}}\right)$
for any reasonable $s$$\qed$
Lemma 6
  1. a)
    The sequence $(p_{2^{k}+1,k+4})_{k=2}^{\infty}$ is descending.
  2. b)
    The sequence $(p_{2^{k}+1,k+5})_{k=3}^{\infty}$ is ascending.
Proof 11
Let $x=2^{k}$ and $t\in\{0,1\}$. In advance we define
$\kappa(k,t):=(-1)^{k+4+t}2^{-\frac{(k+4+t)(k+3+t)}{2}}r_{k+4+t}2^{-2x-1}$
and
$\tau(k,t):=[\![2\!\!\not|(k+t)]\!](-1)^{k+t+3}2^{-\frac{(k+t+3)(k+t+2)}{2}}2r_{k+t+3}\left(\left(\frac{3}{4}\right)^{2x+1}-\left(\frac{1}{2}\right)^{x+2}\right)~{},$
where $[\![\mathrm{cond}]\!]$ is the Iverson bracket of the condition $\mathrm{cond}$. Realize that for $t\in\{0,1\}$ and $k\geqslant 5$, $|\tau(k,t)+\kappa(k,t)|<2^{-50}<10^{-15}$. Now, consider the differences between the consecutive elements of sequences:
[…]
$+$ $\sum_{i=0}^{k+3+t}(-1)^{i}2^{-\frac{i(i-1)}{2}}r_{i}r_{k+t+4-i}\left[\left(1-\frac{2^{-5-t+i}}{x}\right)^{2x+1}-\left(1-\frac{2^{-4-t+i}}{x}\right)^{x+2}\right]$
$=$ $\sum_{i=0}^{\left\lfloor\frac{k+2+t}{2}\right\rfloor}\left\{2^{-i(2i-1)}r_{2i}r_{k+t+4-2i}\left[a(2^{5+t-2i}x)^{2x+1}-a(2^{4+t-2i}x)^{x+2}\right]\right.$
-- $\left.2^{-(2i+1)i}r_{2i+1}r_{k+t+3-2i}\left[a(2^{4+t-2i}x)^{2x+1}-a(2^{3+t-2i}x)^{x+2}\right]\right\}+(\tau+\kappa)(k,t)$
$=$ $\sum_{i=0}^{\left\lfloor\frac{k+t+2}{2}\right\rfloor}2^{-i(2i-1)}r_{2i+1}r_{k+t+4-2i}\left[a(2^{2i+1})\left(a(2^{5+t-2i}x)^{2x+1}-a(2^{4+t-2i}x)^{x+2}\right)\right.$
-- $\left.2^{-2i}a(2^{4+t-2i}x)\left(a(2^{4+t-2i}x)^{2x+1}-a(2^{3+t-2i}x)^{x+2}\right)\right]+\tau(k,t)+\kappa(k,t)~{}.$
Let us define $u_{t}:=2^{5+t-2i}$ (note that $u_{t}$ depends on $i$, but we abbreviate the notation for conciseness) and
$W_{t}(i)$ $:=a(2^{2i+1})\left(a(u_{t}x)^{2x+1}-a\left(\frac{u_{t}}{2}x\right)^{x+2}\right)$
$-2^{-2i}a\left(\frac{u_{t}}{2}x\right)\left(a\left(\frac{u_{t}}{2}x\right)^{2x+1}-a\left(\frac{u_{t}}{4}x\right)^{x+2}\right)$
and consider an upper bound of the last term:
[…]
$-2^{-2i}\left(a\left(\frac{u_{t}}{4}x\right)^{x}\left(a\left(\frac{u_{t}}{4}x\right)+\frac{x+1}{\frac{u_{t}^{2}}{4}x^{2}}\right)-a\left(\frac{u_{t}}{2}x\right)a\left(\frac{u_{t}}{4}x\right)^{x+2}\right)$
=$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})a(u_{t}x)^{2x+1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{u_{t}^{3}x}-\frac{1}{u_{t}^{3}x^{2}}\right)$
$-2^{-2i}a\left(\frac{u_{t}}{4}x\right)^{x}\frac{1}{x}\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{28}{xu_{t}^{2}}+\frac{32}{x^{2}u_{t}^{3}}\right)~{}.$
Note that $2i\leqslant k+2+t$, so ${8}/{x}\leqslant u_{t}$ and in consequence:
  • $06u_{t}-{28}/{x}>{20}/{x}>$ 0,
  • $u_{t}\cdot(3u_{t}+1)-\frac{u_{t}+1}{x}-\frac{1}{x^{2}}$ $\geqslant u_{t}\cdot\left(\frac{24}{x}+1\right)-\frac{u_{t}+1}{x}-\frac{1}{x^{2}}$
    $\geqslant\frac{8}{x}-\frac{1}{x}+\frac{23u_{t}}{x}-\frac{1}{x^{2}}\geqslant\frac{7}{x}+\frac{183}{x^{2}}>0~{}.$
Hence
[…]
$-2^{-2i}a\left(\frac{u_{t}}{4}x\right)\exp\left(-\frac{4}{u_{t}}\right)D(x;2i-3-t)\frac{1}{x}\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{28}{xu_{t}^{2}}+\frac{32}{x^{2}u_{t}^{3}}\right)$
=$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})\exp\left(-\frac{2}{u_{t}}\right)\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{xu_{t}^{3}}-\frac{1}{x^{2}u_{t}^{3}}\right)$ (12)
$-\frac{2^{-2i}}{x}\!\exp\left(-\frac{4}{u_{t}}\right)\!\!\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{32+48u_{t}+28u_{t}^{2}}{u_{t}^{4}x}-\frac{128+64u_{t}-\frac{703}{2}u_{t}^{2}+\frac{259}{4}u_{t}^{3}}{u_{t}^{6}x^{2}}\right.$
$+\left.\frac{512+128u_{t}-1150u_{t}^{2}+\frac{909}{2}u_{t}^{3}}{u_{t}^{7}x^{3}}-\frac{4608-1024u_{t}+530u_{t}^{2}}{u_{t}^{7}x^{4}}+\frac{4096+16u_{t}^{2}}{u_{t}^{8}x^{5}}\right)$
We denote the upper bound (12) by $U_{t}(x;u_{t}(i))$.
Analogically we would like to establish a lower bound of $W_{t}(i)$:
[…]
-- $2^{-2i}a\left(\frac{u_{t}}{2}x\right)\!\left(a\left(\frac{u_{t}}{2}x\right)^{2x+1}\!-a\left(\frac{u_{t}}{2}x\right)^{2x+2}\left(a\left(\frac{u_{t}}{2}x\right)^{2}-\frac{x+2}{\frac{u_{t}^{2}}{4}x^{2}}\right)\right)$
=$\stackrel{{\scriptstyle W}}{{=}}$ $a(2^{2i+1})a\left(\frac{u_{t}}{2}x\right)^{x-1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{10u_{t}+1}{u_{t}^{3}x}+\frac{8}{u_{t}^{3}x^{2}}\right)$ (13)
-- $2^{-2i}a\left(\frac{u_{t}}{2}x\right)^{2x+2}\frac{1}{x}\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{4u_{t}+8}{u_{t}^{3}x}-\frac{8}{u_{t}^{3}x^{2}}\right)$
Now from ${8}/{x}\leqslant u_{t}$ we attain
$u_{t}\cdot(3u_{t}+1)-\frac{10u_{t}+1}{x}$ $=u_{t}\cdot\left(\frac{7u_{t}}{4}+\frac{7}{8}\right)+(10u_{t}+1)\left(\frac{u_{t}}{8}-\frac{1}{x}\right)$
$\geqslant u_{t}\cdot\left(\frac{7u_{t}}{4}+\frac{7}{8}\right)>0$ (14)
and
$u_{t}\cdot(6u_{t}+4)-\frac{4u_{t}+8}{x}-\frac{8}{x^{2}}\geqslant\frac{48u_{t}}{x}+\frac{32}{x}-\frac{4u_{t}+8}{x}-\frac{8}{x^{2}}\geqslant\frac{24}{x}+\frac{344}{x^{2}}>0~{}.$
Hence
[…]
$-2^{-2i}\exp\left(-\frac{4}{u_{t}}\right)a\left(\frac{u_{t}}{2}x\right)\frac{1}{x}\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{4u_{t}+8}{u_{t}^{3}x}-\frac{8}{u_{t}^{3}x^{2}}\right)$
=$\stackrel{{\scriptstyle W}}{{=}}\frac{a(2^{2i+1})}{x}\exp\left(-\frac{2}{u_{t}}\right)\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{2+5u_{t}+4u_{t}^{2}}{u_{t}^{4}x}-\frac{2+4u_{t}-\frac{575}{32}u_{t}^{2}+\frac{387}{32}u_{t}^{3}}{u_{t}^{6}x^{2}}\right.$
$\left.+\frac{2+20u_{t}-\frac{511}{32}u_{t}^{2}+\frac{261}{16}u_{t}^{3}}{u_{t}^{7}x^{3}}-\frac{16+\frac{1}{4}u_{t}^{2}}{u_{t}^{7}x^{4}}\right)$ (15)
$-2^{-2i}\exp\left(-\frac{4}{u_{t}}\right)\frac{1}{x}\left(\frac{6u_{t}+4}{u_{t}^{2}}-\frac{16u_{t}+16}{u_{t}^{3}x}+\frac{16}{u_{t}^{4}x^{2}}+\frac{16}{u_{t}^{4}x^{3}}\right)~{}.$
Denote the lower bound (15) by $L_{t}(x;u_{t}(i))$.
Now we show that $W_{t}(i)>0$ for $i\geqslant 1$. Indeed, after reducing the redundant terms from inequality (13), together with inequality (14), we can obtain
$W_{t}(i)>\frac{a\left(\frac{u_{t}}{2}x\right)^{x}}{xu_{t}^{2}}\left(a(2^{2i+1})\frac{14u_{t}+7}{8}-2^{-2i}(6u_{t}+4)\right)$ (16)
Let us denote $C(t,x):=\frac{a\left(\frac{u_{t}}{2}x\right)^{x}}{xu_{t}^{2}}$. In case $i\geqslant 2$, we may attain
(16$)\geqslant\frac{C(t,x)}{256}\left(31(14u_{t}+7)-96u_{t}-64\right)>0$. When $i=1$, then $u_{t}\geqslant 8$, so (16$)\geqslant\frac{C(t,x)}{64}\left(7(14u_{t}+7)-96u_{t}-64\right)=\frac{2u_{t}-15}{64}\geqslant\frac{1}{64}$.
Thanks to the property $W_{t}(i)>0$ for $i\geqslant 1$, we may subtly neutralize the influence of $r_{k+5-i}$ in the considered sum:
$\sum_{i=0}^{\left\lfloor\frac{k+2}{2}\right\rfloor}2^{-i(2i-1)}r_{2i+1}r_{k+5-i}W_{0}(i)<r_{k+5}\sum_{i=0}^{\left\lfloor\frac{k+2}{2}\right\rfloor}2^{-i(2i-1)}r_{2i+1}W_{0}(i)~{}.$
Naturally we may consider $U_{0}(x;u_{0}(i))$ instead of $W_{0}(i)$ numerically for $i\leqslant 4$:
$\sum_{i=0}^{4}2^{-i(2i-1)}r_{2i+1}U_{0}(x;u_{0}(i))\stackrel{{\scriptstyle W}}{{=}}-8.294491525704523\ldots\cdot 10^{-6}+\frac{0.15588\ldots}{x}\\ +\frac{0.00407163\ldots}{x^{2}}-\frac{0.0298032\ldots}{x^{3}}+\frac{0.0198815\ldots}{x^{4}}-\frac{0.00785419\ldots}{x^{5}}~{},$
so for $x\geqslant 2^{15}\;(k\geqslant 15)$, $\sum\limits_{i=0}^{4}2^{-i(2i-1)}r_{2i+1}W_{0}(i)\leqslant-3.53741\cdot 10^{-6}$. Moreover we may bound $W_{0}(i)$ by $a(2^{5-2i}x)^{2x+1}$ for the rest of the sum:
[…]
so $p_{2^{k+1}+1,k+5}-p_{2^{k}+1,k+4}<0$ for $k\geqslant 15$.
However, according to Theorem 4, we also present the numerical values of the sequence $(p_{2^{k}+1,k+4})_{k=2}^{14}$ in the Table 2. We can now easily see that for any $k\geqslant 2$ we attained $p_{2^{k+1}+1,k+5}-p_{2^{k}+1,k+4}<0$
k $p_{2^{k}+1,k+4}$ k $p_{2^{k}+1,k+4}$ k $p_{2^{k}+1,k+4}$
2 $0.0000305176\ldots$ 7 $0.0000189841\ldots$ 12 $0.0000185484\ldots$
3 $0.0000256707\ldots$ 8 $0.0000187590\ldots$ 13 $0.0000185413\ldots$
4 $0.0000221583\ldots$ 9 $0.0000186466\ldots$ 14 $0.0000185378\ldots$
5 $0.0000203424\ldots$ 10 $0.0000185904\ldots$
6 $0.0000194356\ldots$ 11 $0.0000185624\ldots$
Table 2: Numerical values of the sequence $(p_{2^{k}+1,k+4})_{k=2}^{14}$.
Moreover, realize that ${r_{k+5}}/{r_{k+3}}<1.1$ for any $k\geqslant 3$, so
$\sum\limits_{i=0}^{1}2^{-i(2i-1)}r_{2i+1}1.1^{1-i}L_{1}(x;u_{1}(i))\stackrel{{\scriptstyle W}}{{=}}0.00128843\ldots+\frac{0.00212699\ldots}{x}$
$-\frac{0.00326251\ldots}{x^{2}}+\frac{0.000219133}{x^{3}}-\frac{3.50924875\ldots\cdot 10^{-7}}{x^{4}}$
For any possible $x\geqslant 8$ ($k\geqslant 3$), $\sum_{i=0}^{1}2^{-i(2i-1)}r_{2i+1}1.1^{1-i}L_{1}(x;u_{t}(i))>0.0015$. We already know that $W_{1}(i)$ are positive for $i>1$, so $p_{2^{k+1}+1,k+6}-p_{2^{k}+1,k+5}>0$ for all $k\geqslant 3$$\qed$
We may use Theorem 4 once again to see that ${p_{2^{6}+1,10}}/{p_{2^{6}+1,11}}=129.454\ldots>2^{7}$ and ${p_{2^{7}+1,11}}/{p_{2^{7}+1,12}}=125.065\ldots<2^{7}$. Together with Lemma 6 we may easily attain Claim 1 and we instantly see that this Claim cannot be extended continuously for $k<7$.
Lemma 7
Let $2\leqslant l\leqslant n$ and assume that $\alpha_{i}=2^{i-2}\frac{p_{n,l-i}}{p_{n,l-i+1}}$ for $i\in[0:2]$ and ${\alpha^{\prime}}_{j}=2^{j-2}\frac{p_{n+1,l-j}}{p_{n+1,l-j+1}}$ for $j\in[0:1]$.
If $0\leqslant\alpha_{2}<\alpha_{1}<\alpha_{0}$, then $0<\alpha^{\prime}_{1}<\alpha^{\prime}_{0}$.
Proof 12
Realize that $p_{n+1,l-i+1}=p_{n,l-i+1}(1-2^{-l+i-1}+2^{-l+2}\alpha_{i})$ for $i\in[0:2]$, so for $j\in[0:1]$,
$\alpha^{\prime}_{j}$ $=\frac{p_{n+1,l-j}}{2^{2-j}p_{n+1,l-j+1}}=\frac{p_{n,l-j}(1-2^{-l+j}+2^{-l+2}\alpha_{j+1})}{2^{2-j}p_{n,l-j+1}(1-2^{-l+j-1}+2^{-l+2}\alpha_{j})}$
$=\frac{\alpha_{j}(1-2^{-l+j}+2^{-l+2}\alpha_{j+1})}{1-2^{-l+j-1}+2^{-l+2}\alpha_{j}}~{}.$
Assume that $\alpha^{\prime}_{1}\geqslant\alpha^{\prime}_{0}$. Then
$A:=\alpha_{1}(1-2^{-l+1}+2^{-l+2}\alpha_{2})(1-2^{-l-1}+2^{-l+2}\alpha_{0})\geqslant\alpha_{0}(1-2^{-l}+2^{-l+2}\alpha_{1})^{2}=:B.$
However, contrary to the assumption,
$A$ $=\alpha_{1}(1-2^{-l+1}+2^{-2l}-2^{-l-1}+2^{-l+2}(\alpha_{0}+\alpha_{2})-2^{-2l+3}\alpha_{0}$
$-2^{-2l+1}\alpha_{2}+2^{-2l+4}\alpha_{0}\alpha_{2})$
$<\alpha_{0}(1-2^{-l+1}+2^{-2l})+\alpha_{1}(2^{-l+2}(2\alpha_{0})+2^{-2l+4}\alpha_{0}\alpha_{1})$
$<\alpha_{0}(1-2^{-l+1}+2^{-2l}+\alpha_{1}(2^{-l+3}+2^{-2l+3}+2^{-2l+4}\alpha_{1}))=B~{}.$
$\qed$
Lemma 8
If for some $n\in\mathbb{N}$, $\eta_{n}=2^{-n}\frac{p_{n,n}}{p_{n,n+1}}$ and
$\eta_{n+1}=2^{-n-1}\frac{p_{n+1,n+1}}{p_{n+1,n+2}}$, then $\eta_{n}<\eta_{n+1}$.
Proof 13
00 0 $=p_{n+1,n+1}-2^{n+1}\eta_{n+1}p_{n+1,n+2}=p_{n,n+1}(1-2^{-n-1})+p_{n,n}2^{-n}$
$-\eta_{n+1}p_{n,n+1}=p_{n,n+1}(1-2^{-n-1}+\eta_{n}-\eta_{n+1})~{},$
but $01-2^{-n-1}>$ 0, so $\eta_{n}<\eta_{n+1}$$\qed$